### v3

# supplementary

Abstract
11footnotetext: Corresponding author: Srijoni Majumdar, School of Computing, University of Leeds, Leeds, UK, E-mail: s.majumdar@leeds.ac.uk
Scaling up deliberative and voting participation is a longstanding endeavor – a cornerstone for direct democracy and legitimate collective choice. Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) unravel new capabilities for AI personal assistants to overcome cognitive bandwidth limitations of humans, providing decision support or even direct representation of human voters at large scale. However, the quality of this representation and what underlying biases manifest when delegating collective decision making to LLMs is an alarming and timely challenge to tackle. By rigorously emulating with high realism more than >50K LLM voting personas in 81 real-world voting elections, we disentangle the nature of different biases in LLMS (GPT 3, GPT 3.5, and Llama2). Complex preferential ballot formats exhibit significant inconsistencies compared to simpler majoritarian elections that show higher consistency. Strikingly though, by demonstrating for the first time in real-world a proportional representation of voters in direct democracy, we are also able to show that fair ballot aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation. This novel underlying relationship proves paramount for democratic resilience in progressives scenarios with low voters turnout and voter fatigue supported by AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer. These interdisciplinary insights provide remarkable foundations for science, policymakers and citizens to develop safeguards and resilience for AI risks in democratic innovations.
1 Introduction
Recent advances in artificial intelligence (AI) provide new unprecedented opportunities for citizens to scale up participation in digital democracy [1, 2, 3]. Generative AI in particular, such as large language models (LLMs), have a potential to overcome human cognitive bandwidth limitations and digitally assist citizens to deliberate and decide about public matters at scale [4, 5, 6, 7, 8]. This is by articulating, summarizing and even providing syntheses of complex opinions [9, 10, 7, 8], with a potential to mitigate for the voter fatigue and reduced voters turnout [11, 10], while fostering common ground for compromises, consensus and lower polarization [11, 12, 10, 13, 4]. However, understanding the implications and risks of using large language models for decision support/recommendations or even direct representation of human voters is a pressing challenge [14, 15, 16, 17]. In this article we study the inconsistencies of large language models to approximate both individual and collective choices as well as the inconsistencies of large language models across different preference elicitation methods. We explore causal links of these inconsistencies to potential cognitive biases triggered by the input to large language models based on which choices are made.
Generative AI voting: a converging advance with inevitable challenges. Large language models have been used to predict election winners using sensitive demographic data about the political profile of individuals [18]. They have also been used to predict pairwise comparisons of proposals for constitutional changes [6] and summarized opinions expressed with free-form text [19]. However, little is known whether this AI predictive capability can expand to more complex collective decision-making processes that involve more options to choose from, with more complex preference elicitation methods [5].
Participatory budgeting [20] is one such process put under scrutiny in this article. Here city authorities distribute a public budget by letting citizens propose their own project ideas, which they vote for and often implement themselves [21]. Projects may be pertinent to different impact areas (e.g. environment, culture, welfare), beneficiaries (e.g. elderly, children) and can have different costs. Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares[22, 23]. Recently though, a reinforcement learning approach has been used to elect winners by assisting voters to reach a consensus [24]. So far, the scope and citizens’ engagement in participatory budgeting campaigns remains to a large extent a one-shot and highly local [25, 11]. What if there was a way to geographically spread such decision-making processes and run them on a regular basis? With such complexity and degree of design freedom, scaling up participatory budgeting turns into the ultimate democratic blueprint to assess capabilities and risks of generative AI voting.
Inconsistencies among different preference elicitation methods, known also as intransitivity of preferences [26], have not been studied so far for large language models. However, there is evidence that such inconsistencies do manifest even in human voters [22, 23, 27], especially under polarized contexts [28] that are usually subject of biases [29, 30]. Whether potential biases that explain the inconsistencies between human and AI choices are of different nature than the ones between different input voting methods is an open question studied in this article.
How resilient representative voting outcomes are with generative AI. We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate (see Figure 1d). In other words, we explore whether the inconsistencies of voting outcomes because of low voting turnout are higher than inconsistencies originated from generative AI representatives of abstained voters.
The inconsistencies of generative AI, along with the potential biases explaining these inconsistencies, are systematically studied here for the first time using a novel factorial design based on real-world empirical evidence. It consists of 5 dimensions (see Figure 1): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [31] as well as the mock and actual participatory budgeting voting of the ‘City Idea’ campaign of 2023 in Aarau, Switzerland [32] are studied. They come with different number of ballot options and voters, see Figure 1a and Section 4.2. (ii) Personal human traits - for each voter, multiple incremental levels of additional information are given as input to large language models to generate ballots. This information includes socio-demographics, political preferences, preferences for project categories and preferences for qualities of voting outcomes (popular vs. fair election of winners, individual vs. collective benefit). These are derived from real-world evidence including survey data collected along the voting campaigns (see Section 4.1, and Table LABEL:table:prologue1 and LABEL:table:prologue2). (iii) Voting input methods - four methods with incremental level of complexity and expressiveness are compared [28, 33]. These include single choice for all voting scenarios, approvals, score (assigning a preference score to each option) and cumulative voting (distributing a number of points over the options) [34, 35, 36] for the participatory budgeting scenarios. (iv) AI models - Three large language models are assessed along with a more mainstream predictive machine learning (ML) model used as a benchmark. GPT 3, GPT 3.5, and Llama2 are chosen covering a large spectrum of capabilities in open and closed generative AI [37], see Section 4.1 for more information. (v) Voting aggregation methods - majority is used for the scenario of US elections. The standard utilitarian greedy and the method of equal shares [38] are used for the participatory budgeting scenarios. The former method iterates to elect the next most popular project as long as the budget is not exhausted. It has been the standard method used in real-world. The latter method elects the projects that make the most people ‘happy’, meaning a combination of projects that maximizes a proportional representational of voters’ preferences. This is the method that was actually used in City Idea campaign to elect the winners [39], providing significance and realism for the findings of this study. See section 4.1 for more information.
The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the colored boxes in Figure 1a. This broad spectrum of analysis based on real-world evidence allows us to generalize the findings of the study and make them highly relevant for a broad spectrum of research communities and policymakers.
Figure 1: An overview of the generative AI framework, along with the factorial design and the biases studied in this article. (A) The factorial design with the 5 studied dimensions: (i) real-world voting scenarios in the context of participatory budgeting and national elections. (ii) Various combinations of persona human traits based on which AI personas are created. (iii) Four voting input methods. (iv) Four AI models, three large language models and a predictive machine learning model (benchmark). (v) Voting aggregation methods for elections and participatory budgeting. The studied combinations for each voting scenario are marked with different colors, see also Table 1. (B) The framework of generative AI voting. For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona. The input is the personal human traits, the voting options and the ballot format with instructions for the voting persona about how to make a choice. This choice is the output of the persona. Both human and AI choices are aggregated using a voting aggregation method. The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies. (C) The personal human traits are mapped to cognitive biases. Section 4.4 illustrates choice inconsistencies and their origin to potential cognitive biases. (D) Hypothesizing for a recovery of consistency in voting outcomes under low voters turnout using AI representatives for abstained human voters.
Assessing generative AI voting in action. Voting personas are constructed using input prompts of large language models as depicted in Figure 1b. This designed process aims to emulate the three voting scenarios with the different settings of Figure 1a. Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study in SI) and an instruction to vote according to the ballot format. The individual and collective real-world choices of humans and AI personas are compared for the first time by measuring the overlap (Jaccard similarity) of their selected/winning options as an indicator of consistency [28]. These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), which are fed in a neural network (see Section 4.5). Based on a systematic mapping of human personal traits to the cognitive biases as illustrated in Figure 1c (see Section 4 for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies. This novel analysis is designed to provide a significant conceptual advance to understand how voting design reinforces or mitigates different AI biases in real-world practice.
2 Results
The following three key results are illustrated in this article:
1. Fair voting methods to elect winners are more resilient to human-AI inconsistencies, demonstrating a striking underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2). These inconsistencies are particularly prominent in complex ballot formats with a large number of alternatives, while simple majoritarian voting tends to be highly consistent. GPT 3.5 shows transitivity of preferences among different ballot formats (Figure 3).
2. By representing half of the human voters who abstain with AI, a significant lost consistency is recovered in voting scenarios with low number of alternatives and fair voting aggregation methods to elect winners (Figure 4). For more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI in single-choice majoritarian voting with two alternatives.
3. Conformity biases explain consistency of human-AI choice, while unconscious biases explain inconsistency (Figure 5). Political/ideological profile determines the human-AI consistency in single-choice majoritarian voting with two candidates. Time discounting factors explain transitivity of AI preferences between different ballot formats, while unconscious biases explain intransitivity.
2.1 Fair collective choice is resilient to human-AI inconsistencies
Figure 2 illustrates the human-AI consistency in individual and collective choices. The consistency of individual choice remains poor in complex ballot formats with several alternatives. On average, it is 4.5% and 27.2% for the actual and mock voting in participatory budgeting, yet it is 82.3% for the binary majoritarian US elections. GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 5.2% and 5.31% higher than GPT 3 and Llama2. On the contrary, the consistency of the collective choice increases by 51.2% in total. Strikingly, the consistency of equal shares is on average 80.8%, which is 32.8% higher than utilitarian greedy. Even when reducing the number winners of equal shares to the 8 of utilitarian greedy, the consistency remains 29.4% higher.
Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however consistency drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting. Strikingly, consistency of collective choice is significantly higher than individual choice, particularly for the fairer voting aggregation rule of equal shares. GPT 3.5 shows the highest choice consistency and Llama2 the lowest among the large language models, which though remain inferior to a predictive machine learning model. The consistency (y-axis) in individual and collective choice is shown for different AI models (x-axis) across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) mock, and (C) the US national elections of 2012, 2016 and 2020. The consistency in the participatory budgeting scenarios is shown for different voting input methods (left and right). The consistency of collective choice with the two voting aggregation methods of utilitarian greedy and equal shares is also shown. For the actual voting of City Idea, the consistency of equal shares is calculated for all winners and 8 winners (as many as greedy) for a fairer comparison.
Compared to large language models, the machine learning model shows 3.6% higher consistency in individual choice and 6.6% higher in collective choice. All consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a. The removal of project preferences shows the highest consistency reduction of 19.7%, while political interest shows the lowest consistency reduction of 3.4%.
Figure 3 illustrates the transitivity of preferences in different large language models by measuring the consistency of individual choice among different pairs of ballot formats. ‘Approval - cumulative’ voting shows the highest consistency, which is 1.8% higher than ‘single choice - approval’ and 3.5% higher than ‘single choice - cumulative’.
Figure 3: Voting personas based on GPT 3.5 show transitivity of the preferences over the different voting input methods in the actual complex voting scenario. However, GPT 3 and Llama2 remain with a low consistency <0.5. The simpler mock voting scenario shows higher transitivity among all language models. The consistency (y-axis) in individual choice among different pairs of ballot formats (x-axis) is shown for different large language models, humans and the two voting scenarios in the participatory budgeting campaign of City Idea: (A) actual vs. (B) mock.
GPT 3.5 shows 21.1% higher consistency than GPT 3 and 20.2% higher than Llama2. While human transitivity approaches 95.2% on average, AI transitivity remains at 70.1% for GPT 3.5. The mock voting with only 5 alternatives achieves on average 65.5% of consistency, while the actual voting with 33 alternatives shows a consistency of 43.4%.
2.2 AI representatives to recover from low voters turnout
Figure 4a illustrates the capability of AI representatives to recover the consistency of voting outcomes lost by low voters turnouts. Strikingly, up to 50% of AI representatives are sufficient for a maximum consistency recovery of 53.3% under voting turnouts lower than 80% when equal shares is used. This significant recovery is despite the very low consistency of human-AI individual choices in complex ballot formats with several options (Figure 2a). The recovery is also confirmed for equal shares with the top-8 winners, which is equal to the number of winners with utilitarian greedy (on average 18.1% lower consistency recovery). However, AI representatives fail to recover consistency with utilitarian greedy as the voting aggregation method.
Figure 4: By representing up to half of the human voters who abstain from voting with AI, a significant lost consistency in the collective choice is recovered in participatory budgeting under fair voting aggregation methods such as equal shares and under low number of alternatives. However, further consistency loss is observed under a large number of alternatives with the standard utilitarian greedy method as the number of AI representatives increases. In contrast, single-choice majoritarian elections with highly consistent human-AI individual choice maximize consistency recovery for the whole spectrum of voters turnout by maximizing the AI representatives for the voters who abstain. Nevertheless, for more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI. The consistency loss in the voting outcome by low voters turnout (x-axis) is emulated by removing different ratios of human voters. A recovery of consistency (y-axis) is hypothesized by different ratios of abstained voters represented by AI: 25%, 50%, 75% and 100%. The consistency recovery values are the means of all three large language models (see Figure LABEL:fig:LLM_factorial for the results of each large language model). The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea are shown with two voting aggregation methods: utilitarian greedy and equal shares, with all winners and the top-8 (as many as greedy) for a fairer comparison. (C) US elections. (D) Chord visualizations summarizing how the interplay between voters turnout and AI representation yields voting outcomes with consistency gain or loss. (E) Visualizations summarizing the mechanism of consistency recovery through AI representatives for the voters who abstain.
Figure 4b shows that the recovery of consistency is more robust for both voting aggregation methods under low number of alternatives. Consistency is recovered with up to 75% of abstained voters represented by AI under a maximum of 60% voters turnout, after which 50% of AI representatives are sufficient for consistency recovery.
Under the voting scenario of the US elections that comes with highly consistent human-AI individual choices, consistency is fully recovered for $\geq$25of abstained voters represented by AI under a minimum of 50% voters turnout. However, below 50% of voters turnout, election results are very sensitive and consistency recovery is guaranteed only for 100% of abstained voters represented by AI. More detailed results on mixed populations of humans and AI representatives, including the consistency loss when human voters abstain are shown in Figure LABEL:fig:mixed_populations-abstaining.
2.3 Biases explaining AI inconsistencies in choice and preference transitivity
To explain the origin of (in)consistencies in human-AI choice and different voting input methods, a prediction model is designed for each voting scenario. The personal human traits, mapped to different cognitive biases, become the independent variables and the consistency the dependent variable. The prediction model is built using a neural network (see Section 4.4). The relative importance of the personal human traits that explain the AI consistency is calculated using shapley additive explanations (see Section 4.5).
In the case of participatory budgeting in Figure 5a and 5b, the preference for environmental projects (potential conformity bias [40, 41, 42]) shows an average importance of 26.5% (p=0.0258) in explaining human-AI choice consistency (25% and 12% higher consistency among environmental supporters, with p=0.038 for the actual voting scenario). Political belief and ideology are strongly correlated with the actual two alternatives in the US elections (right/conservative for republicans and left/liberal for democrats), therefore, they strongly explain the consistency of human-AI choice (mean relative importance of 98.7%, p<0.02 and 4.8%, p<0.03 respectively). However, this is not the case in participatory budgeting as it is political belief that rather explains inconsistency (relative importance of -51.2% with p<0.02). This demonstrates that the underlying participatory budgeting choices remain politically neutral to a high extent, despite the fact that the City Idea campaign mainly attracted left-wing voters (42.5% versus 31.5% of right wing voters). In contrast, race and discussion of politics explain inconsistencies in the US elections with a mean relative importance of -48.1% (p<0.04) and -64.7% (p<0.04) respectively. Inconsistency is 33% higher than consistency among white people (p<0.019), while for people who discuss politics inconsistency is 63.2% higher that the ones who do not (p=0.046).
Figure 5: Conformity bias, such as preference for environmental projects, positively contribute to the consistency of human-AI choice. On the other hand, unconscious biases such as race and discussion of politics cause inconsistencies. For US elections though, political belief and ideology strongly contribute to consistency. Time discounting factors such as preference for sport, and welfare projects as well as affect heuristics such as preference for projects that benefit families positively contribute to the consistency of large language models among the voting input methods. The relative importance (x-axis) of the personal human traits (y-axis) to predict the consistency of choice and consistency between voting input methods (single choice vs. cumulative/score) is shown for different large language models and the machine learning model. This is calculated using Shapley Additive Explanations. The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea is shown along with the (C) US elections. The ‘*’ denotes that the feature is statistically significant (p<0.05) for each of the AI models of Llama2, GPT 3, GPT 3.5 and ML respectively. The chord visualizations summarize the biases that explain the (in)consistencies of (D) human-AI choice and (E) voting input method.
The consistency among the voting input methods is explained by preference towards time discounting factors [41] in participatory budgeting, such as sport and welfare projects. The average relative importance is 25.6% and 17.5% respectively (all with p<0.03), with 48.53%, and 37.5% of higher consistency among supporters in the actual voting (p<0.044). Affect heuristics such as preference for projects that benefit families with children also strongly contributes to the consistency with 58.9% higher consistency among supporters in the actual voting (p<0.032). Figures, LABEL:fig:merged3aLABEL:fig:explaination-top-choice, and Table LABEL:table:error illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.
3 Discussion
Generative AI voting: an inevitable race to safeguard democracy. Generative AI voting is likely to emerge as an inevitable technological convergence of AI and e-voting solutions currently adopted in real-world. On the one hand, generative AI and large language models are expected to become more open, pervasive and accessible to citizens [43, 15]. AI personal assistants are already part of everyday life [44, 45, 46], with their generative version expected to follow. On the other hand, the mandate of more direct, secure and active participation in decision making for public matters is expected to further scale up electronic voting solutions and digital platforms. For instance, participatory budgeting elections are mainly conducted digitally, while Estonia has already institutionalized a digital identity for 99% of its citizens as well as electronic voting since 2005 [47]. As Toomas H. Ilves, former president of Estonia, emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen". In the light of these converging advancements, the inter-operation of a generative personal voting assistant with a digital voting platform becomes technologically feasible along with the citizens’ need to have a more direct saying in several public matters and consultations. Therefore, the findings of this study become spot on to understand the implications of such a future, while they are significant to prepare safeguards for digital democracy.
Fair voting design as a safeguard for generative AI voting. Large language models prove not ready yet to accurately represent humans in complex voting scenarios such as participatory budgeting. However, a striking finding of this study, built upon our demonstrated real-world evidence in Aarau, is that consistency of large language models becomes significant under fair voting aggregation methods that promote a proportional representation of voters’ preferences. This motivates a huge opportunity to get democracy "right" in the digital era of AI: move to alternative voting methods that yield fairer voting outcomes for all, while shielding democratic outcomes from AI biases and inconsistencies. The momentum for this win-win paradigm shift is tremendous, which would not be possible otherwise to demonstrate without the real-world empirical evidence based on which this study unfolds: fair voting rules such as the method of equal shares promotes under-represented projects and yield more winning projects with higher proportional representation of citizens’ preferences [39, 22]. Nevertheless, how to scale up these democratic blueprints remains an open question, but AI can play a key role here.
Why fair collective choice is resilient to AI biases and inconsistencies. We provide an explanation of this significant and striking finding. There is evidence that the method of equal shares has an inherent stability in the resulting voting outcomes [22]. For instance, even with 25-50% of the voters, the elected projects do not significantly change [22]. This is despite the fact that projects may end with very different number of total votes, as low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters. Such projects are expected to be an origin of consistency. Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abstaining. As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study becomes the first of its kind and comes with compelling realism and merit for the validity of our significant findings. This comes in stark contrast to other earlier studies [22, 39] that hypothesize the application of equal shares over proposed projects and ballots aggregated with the standard method of utilitarian greedy.
Mitigating low voters turnout: AI representation is not enough, voting design matters. This perspective is assessed for the first time in this study by emulating voting scenarios with varying voters turnout, while letting voters who abstain to have a representation by generative AI to restore the lost consistency of abstaining. Nevertheless, the inherent inconsistencies of generative AI can hinder the effectiveness of AI representatives. Therefore understanding these opposing dynamics comes with significant merit to mitigate the impact of AI biases and inconsistencies in collective decision making. This is by showing striking evidence that voting design has a determining role on how much consistency can be recovered and what effect this can have for the legitimacy of the voting process and outcomes. Recovering the lost consistency under very low voting turnouts in single-choice majoritarian elections requires populations with very high majorities of AI representatives, which turns out to be illegitimate. In contrast, the lost consistency under higher voting turnouts can be restored even with low proportions of AI representatives. The fair method of equal shares is also here a prominent accelerator for the consistency recovery: fewer AI representatives are sufficient to restore consistency even under very low voter turnouts. It is also evident that such consistency recovery could not be achieved with the standard voting aggregation method.
Trustworthy generative AI voting: a call for research and policy action. What information large language models use to reason about voting decisions is influential for different types biases to manifest. Unconscious and conformity biases, related to political belief, race and environment reinforce human choices or yield inconsistencies. Time discounting biases, related to the support of projects on welfare and sport, explain preference transitivity among different ballot formats. These significant findings demonstrate that training data in generative AI voting are expected to be highly impactful for the representation of the voting outcomes in the population. Ethical and democratic guidelines are urgently needed, particularly for the use of (generative) AI in voting processes. For instance, who shall determine the input training data of AI representatives? Should the training data involve only self-determined personal information of voters, or shall these be augmented with more universal knowledge and experts’ opinions? How to protect the privacy and autonomy of voters when training such AI representatives [46]? Will citizens retain power to control AI representatives that reflect their values and beliefs, while remaining accountable? These are some key questions as a basis of a call for action on research and policy in an emerging era of generative AI voting.
4 Methods
We show here how AI representatives are emulated and the real-world data based on which the voting scenarios are constructed. We also illustrate the evaluation approach and the studied human cognitive biases. Finally, the approach to explain the inconsistencies and biases of generative AI is outlined.
4.1 Emulating AI representatives
We introduce a generalized design to create AI representatives using three large language models and a predictive AI method as a benchmark (referred to as ML) to generate choices (ballots) and emulate AI voting representation.
Input to ballot generation. This comprises of (i) defining the voting scenario, the voting alternatives and the personal human traits of the voter as the input context (LLMs) or features (ML) to inform the ballot choice, (ii) generating the ballots and (iii) aggregating the ballots into a collective decision (see Figure 1).
  • (i)
    Personal human traits: Voter’s information is used to define an AI persona of a human voter, i.e. the AI representative (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2 in SI). This includes (i) socio-demographic characteristics (e.g., gender, age, education, household size), (ii) political interests (e.g., ideological profile, political belief), (iii) personal attitude towards project preferences (e.g., prioritization for green initiatives, sustainable transport, elderly care facilities) and (iv) preferences for the qualities of voting outcomes (e.g., favoring cost-effective winning projects vs. popular ones vs. winning projects with proportional representation of citizens’ preferences). These traits are extracted from feedback surveys of voters, who also cast ballots in the voting scenarios and therefore the survey data are linked to the actual voting behavior (see field study design in Section 4.2).
  • (ii)
    Voting alternatives: The three real-world voting scenarios determine the voting alternatives given as input to the AI representatives. They are illustrated in detail in Section 4.2. The actual City Idea scenario comes with 33 project alternatives and their cost, while the mock scenario comes with 5 costed project alternatives. The US elections comes with two alternatives. These three scenarios provide a broad spectrum of decisions spaces to assess the AI representatives.
  • (iii)
    Ballot formats: The two participatory budgeting scenarios are either approvals (n-approval ballot where ‘n’ is the number of projects approved) or cardinal preferences (score or cumulative) [48, 49]. In the case of score ballots, the voters provide a score to a project from a specified range such as 1 to 5 with 1 being the lowest preference and 5 being the highest. In the case of cumulative ballots, the voters distribute a number of given points among the projects. Single-choice ballots are derived from single approvals, top scores or top points distributed to a project.
Generative AI. We use proprietary models such as GPT 3.5, GPT3 [50] and the open-source Llama2-13B [51]. GPT 3.5 and GPT 4 have similar performance for these tasks [52], hence we resort to GPT 3.5 for the analysis. Furthermore, both GPT 3.5 and GPT 4 use the “Let’s think step by step” to evaluate a prompt and find a solution that mimics human reasoning [52]. For each of the large language models, predictions are made based on the five combinations of personal human traits (see Figure 1), however the predictions with all models have the highest consistency to human choices, which we use for the analysis.
Predictive AI benchmark. The personal human traits are set as features to predict the ballots using classification methods [53]. We sample 40% of the data and use synthetic minority oversampling [54] to augment the data. This 40% is decided empirically based on the performance of the neural network models. We also divide the prediction into multiple sub-problems, which predict the approvals or assigned scores for every option and then compile the final ballot. This model is then used to predict 60% of the unknown ballots and 40% of the known ballots.
Ballot aggregation. For the US elections, majority determines the winner. For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method or the method of equal shares. Utilitarian greedy simply selects the next most popular project, the one with the most received votes, as long as the available budget is not exhausted. Equal shares provides a proportional representation of voters’ preferences by dividing the budget equally among voters as endowments. This part of the budget can only be used to fund projects that the voter has voted for. The method goes through all project options, beginning from the projects with the highest number of votes. It selects a project if it can be funded using the budget shares of those who voted for the project. A full explanation of the is out of the scope of this article and can be found in earlier work [38, 48, 49]. In practice, the method of equal shares is likely to sacrifice an expensive popular project to select instead several low-cost projects, which, in combination, satisfy the preferences of more voters [39, 22]. As a result, equal shares tends to elect more winning projects. Because of this effect, it likely that consistency measurements based on the Jaccard similarity yield higher values for equal shares. This is the reason we control for the number of winning projects in equal shares by counting a subset of the most popular winning projects which is equal in number with the winners of the utilitarian greedy method.
Emulated voting processes with human voters and AI representatives. We emulate the following voting processes: (i) Human voters exclusively with 100% turnout. (ii) Human voters exclusively with varying turnout in the range [10%,100%] with a step of 10%. (iii) Mixed populations of human voters and AI representatives assessed of whether they can recover the consistency of voting outcomes lost by different human voters turnouts in the range [10%,100%] with a step of 10%. The ratio of AI representatives for the abstained voters varies in the range [25%,100%] with a step of 25%. (iv) AI representatives exclusively.
Each scenario of voters turnout is repeated 10 times by selecting a random subset of voters who abstain. For each turnout level, the AI representation of abstained voters is also repeated 10 times by selecting a random subset of the abstained voters to represent by AI.
Data collection infrastructure. Generative AI choices are collected via API prompts to large language models from 16th of June 2023 to 8th of November 2023. We prompted the large language models using the zero-shot learning feature [50], which does not require any specific fine-tuning. We use chain of thought prompting [55] to provide a comprehensive and systematic flow of information for better interpretability. The large language models run with temperature settings of 0.2 to 0 (20 runs for each temperature setting), for a focused decision with less randomness [56]. As we are dealing with a significantly large decision space, particularly the 33 projects in the actual voting, running with a high-temperature setting can lead to random decision-making. To maintain consistency, we retain the projects that are approved at the deterministic temperature setting (the responses are easily reproducible at temperature setting = 0). Furthermore, if a project is selected both in 0.1 and 0.2, we include the approval for that project in the ballot to also incorporate diversity. In the case of score voting, we take the frequency of the distinct scores assigned to a project across different temperature setting and add the score with the highest frequency. However, for the actual City Idea scenario for which we emulate cumulative voting, the total score provided in a ballot should be lower than the total number of points. Therefore, it would not be be accurate to derive a score for a project from multiple outputs corresponding to different temperature settings, hence we conducted experiments only with the temperature setting of 0.2. The decoding algorithm for the large language model is set to the nucleus sampling approach [55]. Compared to purely random sampling, it avoids generating outliers or improbable sequences of choices.
For the predictive AI choices, we use neural networks with 3 hidden layers and the relu [57] activation function for the sub-networks in the neural network. We employ the ensemble method [58] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [59]. This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data.
4.2 Voting datasets
Table 1 outlines the settings of the emulated voting scenarios.
US elections. The 2012, 2016, and 2020 survey waves of the American National Election Study (ANES) [60] are used. They encompass 4,500 votes together with the respective voter information for each of these three years. This mainly includes socio-demographic characteristics: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence. A total of 9 elections are emulated using Llama, GPT-3.5 and the predictive ML model based on single-choice ballots, with winners decided by majority for the three years. We extracted information from 20,650 voters of the ANES dataset [60] and after pre-processing (removing voters with missing values in more than 5 of the personal traits), a set of 17,010 voters is used. The examples of the prompts used to generate the AI choices can be found in Table LABEL:table:Anes.
City Idea: Participatory budgeting campaign in Aarau, Switzerland. The data from a recent innovative participatory budgeting field study are used [32, 39], which was conducted with ethical approval from University of Fribourg (#2021-680). It run in 2023 and is rigorously designed to assess the application of equal shares for the first time in real world, in combination with cumulative voting, using the open-source Stanford Participatory Budgeting platform [33]. As such, the field study includes a survey conducted before voting linking the choices of survey respondents and voters. We use the following personal human traits from the survey (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2): (i) 9 key socio-demographic characteristics (e.g., age, citizenship, education) and 2 political interests (political beliefs and trust in democracy), (ii) preferences for 9 different types of projects and 6 beneficiaries and (iii) 4 types of preferences for qualities of the voting outcome.
Table 1: The studied dimensions across three real-world voting scenarios. They provide the necessary diversity to generalize the findings of this study as they include different scales in number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases including both generative and predictive AI methods.
Studied factors US Elections City Idea City Idea
2012, 2016, 2020 [Mock] [Actual]
Ballot Input
Personal human traits
Socio-demographics
Preference for voting outcome qualities
Project preferences
Political interests
Ballot format
Single-choice
Approval
Score
Cumulative
Alternatives for voting 2 5 33
Ballot Generation
Generative AI
GPT 3
GPT 3.5
Llama2-13B
Predictive AI
Neural Networks
Ballot Aggregation
Majority
Utilitarian greedy
Equal shares
Voters $\sim$17,000 (across 3 years) 3,134 505
Emulated elections 9 91 27
Two participatory budgeting voting scenarios are studied in the context of the real-world campaign of City Idea:
  • (i)
    Mock Voting: Five hypothetical costed projects belonging in different categories are put for mock voting as part of the initial survey. Table LABEL:table:prologueS illustrates the project alternatives and their cost. The choice of 3134 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting. This allows us to emulate a total of 45 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits + 1 for the predictive ML model.
  • (ii)
    Actual Voting: Using the Stanford Participatory Budgeting platform [33], 1703 voters cast their vote using cumulative ballots by distributing 10 points to at least 3 projects of their preference, out of 33 projects in total (see Table LABEL:table:prologueA for project descriptions). A subset of 505 of these voters, which participated in the initial survey and provided their personal human traits, are used to construct the AI representatives. The ballot formats of single choice and approvals are derived from the cumulative ballots by taking the project with the most points and the projects that received any point respectively. This allows us to emulate a total of 91 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits.
The examples of the prompts used to generate the AI choices in both mock and actual voting can be found in Table LABEL:table:prologuePersona.
4.3 Evaluation of choices by AI representatives
The emulated elections with AI representatives are compared to the real-world elections of human voters at two levels: (i) individual choice, i.e. the ballots, and (ii) collective choice, i.e. the resulting voting outcomes. Consistency is the main measure of assessment.
Consistency of individual choice. Consistency is assessed at two levels:
  • (i)
    Human-AI consistency of individual choice: This is the Jaccard similarity [61] of individual choices made by AI representatives with the original choices of the humans they represent. For single-choice ballots, this is either 1 (consistent) or 0 (inconsistent) for each voter and the respective AI representative. For approval voting, the AI and human choices are binary sequences representing the ballot, containing 1 (approved) or 0 (not approved) for each alternative. The Jaccard similarity between these sequences is the ratio between the number of alternatives approved by both the human and the AI representative over the number of alternatives approved by at least one of them. For score and cumulative voting, projects are ranked based on the scores and points received respectively. Each top-k project of a human voter is compared to the respective top-k project by its AI representative. Matching is indicated by 1 (consistent), otherwise 0 (inconsistent) indicates no matching. The mean number of matchings among all ranked projects measures the consistency for these ballot formats.
  • (ii)
[…]
Consistency of collective choice. It is defined for two voting processes that elect winners from the same alternatives, but have different voters (human, AI representatives, mixed), and, as a result, can yield different voting outcomes. The voting outcomes with the winners are first calculated using one of the ballot aggregation methods, i.e. majority, utilitarian greedy or equal shares. These outcomes are represented as binary sequences containing either 0 (not winner) or 1 (winner) for each alternative. Consistency is calculated using the Jaccard similarity [61] between these binary sequences.
Consistency recovery in collective choice with AI representatives. It is determined here for voting processes with varying voters turnout, in which abstained voters result in consistency loss, which can be recovered if a portion of these abstained voters are represented by AI. It is measured as follows:
$\frac{consistency\ [\text{human\ voters - abstained\ voters + AI\ representatives}]-consistency\ [\text{human\ voters - abstained\ voters}]}{1-consistency\ [\text{human\ voters - abstained\ voters}]},$
where the voters turnout human voters - abstained votershuman voters + abstained $\frac{\text{human\ voters - abstained\ voters}}{\text{human\ voters + abstained\ voters}}$ varies in the range [10%,90%] with a step of 10%, and AI representation AI representativesabstained $\frac{\text{AI\ representatives}}{\text{abstained\ voters}}$ varies in the range [25%,100%] with a step of 25%.
4.4 Human cognitive biases in AI collective decision making
Human choices are significantly influenced by potential cognitive biases that are often a manifestation of socio-economic characteristics, conditions of life quality, satisfaction with the available public amenities and the overall life experiences of an individual [41]. Here we map self-reported personal human traits with which large language models are prompt to potential underlying human cognitive biases, as a way to explore whether they are reinforced by the language models and become more likely to manifest in the AI choices. Figure 1c outlines the mapping we study based on a systematic and comprehensive review of relevant literature. The following types of biases are determined:
Time-discounting biases. These are characterized by the tendency to receive immediate gratification over a larger but future reward. Projects related to the public space, culture, or welfare are often one-off projects, such as an annual festival, a cinema night, or educating asylum children (alternatives proposed for the participatory budgeting campaign in Aarau [32]), where expected benefits are experienced in the short term. Therefore, such project are subject of time-discounting biases [41]. In contrast, the construction of a road, for example, incurs investment costs several years before the actual project implementation. Such investments create a sustainable reform that benefits all in the long run. Even though 72% of the public transportation infrastructure in European cities ends up in cost overruns, people still vote for these projects due to an underlying necessary optimism to improve transport [41, 39]. We refer to this economic discounting over a long-term sustainability impact as ‘optimism’ [62, 63].
Surrogation biases. This reflects how humans favor simpler measures to assess the impact over ones that are more precise and harder to evaluate. Korteling et al. [41] argue that these biases manifest when deciding projects with large societal impact such as health or education, while their outcome is subject of different satisfaction levels among citizens. The project outcomes may be perceived as successful by part of society using easy-to-evaluate metrics instead of looking at long-term effects on the community. For instance, a timely vaccination drive may be preferred over significant changes to vaccination protocols or health insurance policies covering vaccination. Hence these projects are likely to be more preferred as they come with more intuitive ways to assess for the broader population. This is reflected by the average winning rate of the health and education-related projects of 38.1% and 36.2% respectively in Poland [39], where participatory campaigns have been actively hosted in the last decade.
Conformity biases. These biases arise out of group pressure under which people make decisions to be socially desirable [40, 42]. It is argued that a conformity bias may induce voting for green alternatives [41]. Green-themed participatory budgeting campaigns have been adopted in European cities such as Lisbon [64], to promote green initiatives, aligning to a culture for more sustainable behavior. Poland runs participatory budgeting campaigns at large scale, which include environment and urban greenery projects. These are within the top-5 most popular projects with an average of 22.5% and 26.5% respectively [39]. Even in Aarau we observe the same trend wherein, environmental friendly projects account for the top-10 most popular projects [32].
Affect heuristic biases. This defines the tendency to make decisions based on what intuitively or emotionally feels right. Affect biases have been studied to analyze the inclusive attitudes most people show towards elderly people [65]. Similar biases also manifest in welfare of children and in inter-generational communication [41, 66]. In Aarau, we observe that 71.3% of the voters prefer projects for younger and elderly people.
Biases for altruism and egotism. Individual interest is often in conflict with the community interest in participatory and collective decision-making processes. While voting, the intrinsic altruism of citizens strongly influences choices. As a result, altruism and egotism are influential for the fairness of voting outcomes and how these outcomes benefit the city in overall [67, 41, 66]. In Aarau, we observe that 67.1% of the voters, who prefer better representation in the outcome are prosocial and prioritize the city-wide benefit (altruism bias) over individual benefit (egotism bias).
Unconscious biases. Human choices are influenced by socio-economic and demographic traits such as race, ethnicity, citizenship, household size and income [68]. Specifically, political ideology and belief shape to a high degree decisions for candidates in elections [69].
4.5 Explainability of generative AI voting
The consistency of the individual AI choices (see Section 4.3) is modeled as the dependent variable in a predictive machine learning framework. Causal relationships that explain how personal human traits (independent variables) contribute to consistency are studied. We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$. Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set. We also assess prediction models with different dependent variables, see Figure LABEL:fig:explaination-top-choice for more details.
Explainability of choices. We introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on the consistency of individual choice. We use model agnostic Shapley Additive Explanations [70] to extract individual contributions of each trait. Results are shown in Figure 5. Feature ablation study [71] is used to calculate the error (loss) in the overall prediction accuracy of the model when a feature is removed (results in Table LABEL:table:error).
Acknowledgements
This work is funded by a UKRI Future Leaders Fellowship (MR/W009560/1): Digitally Assisted Collective Governance of Smart City Commons–ARTIO’. The participatory budgeting data were earlier collected in the context of the City Idea campaign with support from Swiss National Science Foundation NRP77 ‘Digital Transformation’ project (#407740_187249): Digital Democracy: Innovations in Decision-making Processes. The authors would like to thank Thomas Wellings, Regula Hänggli and Dirk Helbing for constructive discussions.
Supplementary Material
Supplementary material is available below.
Author contributions statement
S.M. wrote the manuscript, collected the data, designed and developed the AI models, and analyzed the data. E.E. edited the manuscript and analyzed the data. E.P. wrote the manuscript, conceived the study, designed the AI models and analyzed the data.
Data Availability
The data for US elections are available in the link https://figshare.com/collections/Generative_AI_Voting_-_ANES/7261288.
Code Availability
The codes are available in the link https://github.com/TDI-Lab/Generative-AI-Voting

### v4

@@
 Scaling up deliberative and voting participation is a longstanding endeavor – a cornerstone for direct democracy and legitimate collective choice.
-Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) unravel new capabilities for AI personal assistants to overcome cognitive bandwidth limitations of humans, providing decision support or even direct representation of human voters at large scale.
+Recent breakthroughs in generative artificial intelligence (AI) and large language models (LLMs) unravel new capabilities for AI personal assistants to overcome cognitive bandwidth limitations of humans, providing decision support or even direct representation of abstained human voters at large scale.
 However, the quality of this representation and what underlying biases manifest when delegating collective decision making to LLMs is an alarming and timely challenge to tackle.
-By rigorously emulating with high realism more than >50K LLM voting personas in 81 real-world voting elections, we disentangle the nature of different biases in LLMS (GPT 3, GPT 3.5, and Llama2).
+By rigorously emulating with high realism more than >50K LLM voting personas in 306 real-world voting elections, we disentangle the nature of different biases in LLMs.
 Complex preferential ballot formats exhibit significant inconsistencies compared to simpler majoritarian elections that show higher consistency.
-Strikingly though, by demonstrating for the first time in real-world a proportional representation of voters in direct democracy, we are also able to show that fair ballot aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation.
-This novel underlying relationship proves paramount for democratic resilience in progressives scenarios with low voters turnout and voter fatigue supported by AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer.
+Strikingly though, fair ballot aggregation methods, such as equal shares, prove to be a win-win: fairer voting outcomes for humans with fairer AI representation, especially for voters who are likely to abstain.
+This novel underlying relationship proves paramount for democratic resilience in progressive scenarios with low voters turnout by voter fatigue supported using AI representatives: abstained voters are mitigated by recovering highly representative voting outcomes that are fairer.
 These interdisciplinary insights provide remarkable foundations for science, policymakers and citizens to develop safeguards and resilience for AI risks in democratic innovations.
@@
 Generative AI in particular, such as large language models (LLMs), have a potential to overcome human cognitive bandwidth limitations and digitally assist citizens to deliberate and decide about public matters at scale [4, 5, 6, 7, 8].
-This is by articulating, summarizing and even providing syntheses of complex opinions [9, 10, 7, 8], with a potential to mitigate for the voter fatigue and reduced voters turnout [11, 10], while fostering common ground for compromises, consensus and lower polarization [11, 12, 10, 13, 4].
-However, understanding the implications and risks of using large language models for decision support/recommendations or even direct representation of human voters is a pressing challenge [14, 15, 16, 17].
-In this article we study the inconsistencies of large language models to approximate both individual and collective choices as well as the inconsistencies of large language models across different preference elicitation methods.
-We explore causal links of these inconsistencies to potential cognitive biases triggered by the input to large language models based on which choices are made.
-Generative AI voting: a converging advance with inevitable challenges.
+This is by articulating, summarizing and even providing syntheses of complex opinions [9, 10, 7], with a potential to mitigate for the voter fatigue and reduced voters turnout [11, 10, 12], while fostering common ground for compromises, consensus and lower polarization [11, 13, 10, 14, 4].
+However, understanding the implications and risks of using large language models for decision support, recommendations or even direct representation of human voters is a pressing challenge [15, 16, 17].
+Unraveling inconsistencies in generative AI voting.
+We disentangle the inconsistencies of large language models in both individual and collective choices.
+In particular, we study three manifestations of choice inconsistency as shown in Figure 1a: (i) inconsistency by under-representation as a result of low human voters turnout, (ii) inconsistency by inaccurate approximation of human choice by AI and (iii) inconsistency by intransitivity [18] of AI choice over different ballot formats.
+Intransitivity even manifests among human voters, especially under polarized contexts [19] that are usually subject of biases [20, 21], therefore, we expect LLM choices to be vulnerable to such inconsistencies.
+Whether potential biases that explain the inconsistencies between human and AI choices are of different nature than the ones between different input voting methods is an open question studied in this article.
+We rigorously measure such inconsistencies with a single universal approach grounded to social choice theory [13, 22].
+It exhaustively characterizes the similarity of the two choices (individual or collective) by counting the relative number of Condorcet pairwise matches, see Section 4.3 for further information.
+We also explore causal links of these inconsistencies to potential cognitive biases triggered by the input to large language models based on which choices are made.
+Figure 1: An overview of the studied generative AI voting framework.
+(A) Three manifestations of choice inconsistency are distinguished, measured using Condorcet pairwise matches: (i) inconsistencies by under-representation as a result of low voters turnout, (ii) inconsistencies by inaccuracy of AI choice to approximate human choice and (iii) inconsistency by intransitivity of AI choices over different ballot formats.
+(B) The factorial design with the 7 studied dimensions: (i) Real-world voting scenarios in the context of participatory budgeting and national elections.
+(ii) Various combinations of personal human traits (features) based on which AI voting personas are created.
+(iii) Four ballot formats.
+(iv) Four AI models, three large language models and a predictive machine learning model (benchmark).
+(v) Ballot aggregation methods for elections and participatory budgeting.
+(vi) The three abstaining models that are based for engagement, digital literacy and trust.
+(vii) Participation modalities ranging from exclusive human participation of varying turnout to mixed populations of humans and AI representatives of abstained voters.
+The studied combinations for each voting scenario are marked with different colors, see also Table 1.
+(C) The framework of generative AI voting.
+For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona.
+The input is the personal human traits, the voting options, and the ballot format, with instructions for the voting persona on how to make a choice.
+This choice is the output of the persona.
+Both human and AI choices are aggregated using a ballot aggregation method.
+The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies.
+(D) The personal human traits are mapped to cognitive biases.
+Section LABEL:sec:bias illustrates the origin of choice inconsistencies to potential cognitive biases.
+Generative AI voting: a converging advance technology with inevitable challenges.
 Large language models have been used to predict election winners using sensitive demographic data about the political profile of individuals [23].
 They have also been used to predict pairwise comparisons of proposals for constitutional changes [6] and summarized opinions expressed with free-form text [24].
-However, little is known whether this AI predictive capability can expand to more complex collective decision-making processes that involve more options to choose from, with more complex preference elicitation methods [5].
+However, little is known whether this AI predictive capability can expand to more complex ballot formats that involve more options to choose from [5].
 Participatory budgeting [25] is one such process put under scrutiny in this article.
@@
 environment, culture, welfare), beneficiaries (e.g.
-elderly, children) and can have different costs.
-Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares[22, 23].
+elderly, children) and can have different costs [27].
+Voters can approve, rank or distribute points over their preferred projects, while winners are elected based on the popularity of the projects (utilitarian greedy) or based on a proportional representation of the voters’ preferences (equal shares or Phragmen’s rule) [28, 29].
 Recently though, a reinforcement learning approach has been used to elect winners by assisting voters to reach a consensus [30].
@@
 With such complexity and degree of design freedom, scaling up participatory budgeting turns into the ultimate democratic blueprint to assess capabilities and risks of generative AI voting.
-Inconsistencies among different preference elicitation methods, known also as intransitivity of preferences [26], have not been studied so far for large language models.
-However, there is evidence that such inconsistencies do manifest even in human voters [22, 23, 27], especially under polarized contexts [28] that are usually subject of biases [29, 30].
-Whether potential biases that explain the inconsistencies between human and AI choices are of different nature than the ones between different input voting methods is an open question studied in this article.
 How resilient representative voting outcomes are with generative AI.
-We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate (see Figure 1d).
+We hypothesize that generative AI representatives of human voters build up resilience for representative voting outcomes in voting processes of low turnout if these human voters would otherwise abstain or would not have capacity to actively participate (see participation modalities in Figure 1b).
 In other words, we explore whether the inconsistencies of voting outcomes because of low voting turnout are higher than inconsistencies originated from generative AI representatives of abstained voters.
+This overall consistency recovery by AI representation defines the resilience of a voting outcome in voting scenarios of low voting turnouts and mixed populations of humans and AI representatives of abstained voters.
+Disentangling the role of voting design in generative AI voting.
 The inconsistencies of generative AI, along with the potential biases explaining these inconsistencies, are systematically studied here for the first time using a novel factorial design based on real-world empirical evidence.
-It consists of 5 dimensions (see Figure 1): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [31] as well as the mock and actual participatory budgeting voting of the ‘City Idea’ campaign of 2023 in Aarau, Switzerland [32] are studied.
-They come with different number of ballot options and voters, see Figure 1a and Section 4.2.
+It consists of 7 dimensions (see Figure 1b): (i) Real-world voting scenarios - datasets from the 2012, 2016 and 2020 US national elections [32] as well as data from the 2023 participatory budgeting campaign of ‘City Idea’ in Aarau, Switzerland [33].
+The latter dataset comes with two voting scenarios, a hypothetical one provided to voters before voting via a survey and the actual voting data.
+These three scenarios cover a profound range of different number of ballot options and voters to experiment with, see Figure 1b and Section 4.2.
 (ii) Personal human traits - for each voter, multiple incremental levels of additional information are given as input to large language models to generate ballots.
@@
 These are derived from real-world evidence including survey data collected along the voting campaigns (see Section 4.1, and Table LABEL:table:prologue1 and LABEL:table:prologue2).
-(iii) Voting input methods - four methods with incremental level of complexity and expressiveness are compared [28, 33].
+(iii) Ballot formats - four methods with incremental level of complexity and expressiveness are compared [19, 34].
 These include single choice for all voting scenarios, approvals, score (assigning a preference score to each option) and cumulative voting (distributing a number of points over the options) [35, 36, 37] for the participatory budgeting scenarios.
 (iv) AI models - Three large language models are assessed along with a more mainstream predictive machine learning (ML) model used as a benchmark.
-GPT 3, GPT 3.5, and Llama2 are chosen covering a large spectrum of capabilities in open and closed generative AI [37], see Section 4.1 for more information.
-(v) Voting aggregation methods - majority is used for the scenario of US elections.
-The standard utilitarian greedy and the method of equal shares [38] are used for the participatory budgeting scenarios.
-The former method iterates to elect the next most popular project as long as the budget is not exhausted.
+GPT 3, GPT 3.5, and Llama3-8B are chosen covering a large spectrum of capabilities in open and closed generative AI [38], see Section 4.1 for more information.
+(v) Ballot aggregation methods - majority is used for the scenario of US elections.
+The standard utilitarian greedy, the method of equal shares [39] and the Phragmen’s sequential rule [40] are used for the participatory budgeting scenarios.
+Utilitarian greedy iterates to elect the next most popular project as long as the budget is not exhausted.
 It has been the standard method used in real-world.
-The latter method elects the projects that make the most people ‘happy’, meaning a combination of projects that maximizes a proportional representational of voters’ preferences.
-This is the method that was actually used in City Idea campaign to elect the winners [39], providing significance and realism for the findings of this study.
+The latter two methods refer to as the fair collective choice based on the principle of proportionality: they elect a combination of projects such that a more proportional representation of voters’ choice is achieved by providing to voters power to decide for an equal part of the available budget [39].
+Equal shares is the method that was actually used in City Idea campaign to elect the winners [27], providing a compelling realism for the findings of this study.
 See section 4.1 for more information.
-The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the colored boxes in Figure 1a.
-This broad spectrum of analysis based on real-world evidence allows us to generalize the findings of the study and make them highly relevant for a broad spectrum of research communities and policymakers.
-Figure 1: An overview of the generative AI framework, along with the factorial design and the biases studied in this article.
-(A) The factorial design with the 5 studied dimensions: (i) real-world voting scenarios in the context of participatory budgeting and national elections.
-(ii) Various combinations of persona human traits based on which AI personas are created.
-(iii) Four voting input methods.
-(iv) Four AI models, three large language models and a predictive machine learning model (benchmark).
-(v) Voting aggregation methods for elections and participatory budgeting.
-The studied combinations for each voting scenario are marked with different colors, see also Table 1.
-(B) The framework of generative AI voting.
-For each voter in the real-world voting scenario, a prompt is given to large language models to construct the voting persona.
-The input is the personal human traits, the voting options and the ballot format with instructions for the voting persona about how to make a choice.
-This choice is the output of the persona.
-Both human and AI choices are aggregated using a voting aggregation method.
-The inconsistencies of individual and collective choices for humans and AI personas are assessed, along with potential biases that explain these inconsistencies.
-(C) The personal human traits are mapped to cognitive biases.
-Section 4.4 illustrates choice inconsistencies and their origin to potential cognitive biases.
-(D) Hypothesizing for a recovery of consistency in voting outcomes under low voters turnout using AI representatives for abstained human voters.
+(v) Abstaining models - the voters with low engagement and trust in political processes, as well as low digital literacy represent the ones who are likely to abstain.
+They are identified using available survey data and grounded to well-established literature [12, 41, 42] (see Section 4.1).
+(vi) Participation modalities - we assess the consistency of voting scenarios with full and low turnouts of human voters, partial / full AI representation of abstained voters and AI representation of the whole human population.
+The dimensions of the factorial design are illustrated in Table 1 and the studied combinations are marked with the colored boxes in Figure 1b.
+This broad spectrum of analysis based on real-world evidence allows us to generalize the findings of the study and make them relevant for a broad spectrum of research communities and policymakers.
 Assessing generative AI voting in action.
-Voting personas are constructed using input prompts of large language models as depicted in Figure 1b.
-This designed process aims to emulate the three voting scenarios with the different settings of Figure 1a.
-Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study in SI) and an instruction to vote according to the ballot format.
-The individual and collective real-world choices of humans and AI personas are compared for the first time by measuring the overlap (Jaccard similarity) of their selected/winning options as an indicator of consistency [28].
-These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), which are fed in a neural network (see Section 4.5).
-Based on a systematic mapping of human personal traits to the cognitive biases as illustrated in Figure 1c (see Section 4 for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies.
-This novel analysis is designed to provide a significant conceptual advance to understand how voting design reinforces or mitigates different AI biases in real-world practice.
+Voting personas are constructed using input prompts of large language models as depicted in Figure 1c.
+This designed process aims to emulate the three voting scenarios with the different settings of Figure 1b.
+Each input prompt consists of a standardized description of the voter’s profile (see Section LABEL:sec:study) and an instruction to vote according to the ballot format.
+The individual and collective real-world choices of humans and AI personas are compared for the first time by measuring the relative number of Condorcet pairwise matches as an indicator of consistency [13, 22, 19].
+These consistency values are then becoming the dependent variable to predict using the personal human traits as independent variables (features), fed in a neural network (see Section 4.4).
+Based on a systematic mapping of human personal traits to cognitive biases as illustrated in Figure 1c (see Section LABEL:sec:bias for more detail), this prediction model causally explains the human traits that contribute to inconsistencies and the potential underlying biases that explain these inconsistencies.
+This novel analysis is designed to provide a significant conceptual advance in understanding how voting design reinforces or mitigates different AI biases in real-world practice.
 2 Results
@@
 1.
-Fair voting methods to elect winners are more resilient to human-AI inconsistencies, demonstrating a striking underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2).
+Fair voting methods to elect winners are more resilient to inconsistencies of AI to accurate estimate human choice, demonstrating a striking underlying win-win relationship: fairer voting outcomes for humans with fairer human representation by AI (Figure 2 and 4).
 These inconsistencies are particularly prominent in complex ballot formats with a large number of alternatives, while simple majoritarian voting tends to be highly consistent.
-GPT 3.5 shows transitivity of preferences among different ballot formats (Figure 3).
+AI intransitivity across ballot formats is higher than the minimal one of humans, with higher impact though on collective choice over a large number of alternatives (Figure 3).
 2.
-By representing half of the human voters who abstain with AI, a significant lost consistency is recovered in voting scenarios with low number of alternatives and fair voting aggregation methods to elect winners (Figure 4).
-For more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI in single-choice majoritarian voting with two alternatives.
+AI representation is more effective over voters who is likely to abstain than an arbitrary voter, in particular under fair collective choice (Figure 4).
+Abstaining voters result in a representation deficit restored by AI, while AI representation over arbitrary voters mainly has a noise-reduction effect or voting outcome.
 3.
-Conformity biases explain consistency of human-AI choice, while unconscious biases explain inconsistency (Figure 5).
-Political/ideological profile determines the human-AI consistency in single-choice majoritarian voting with two candidates.
-Time discounting factors explain transitivity of AI preferences between different ballot formats, while unconscious biases explain intransitivity.
+Features of abstaining voters related to their low engagement, digital literacy and trust explain the consistency of their AI representation and transitivity of ballot formats (Figure 5).
+Affect and unconscious biases explain the (in)consistency of human-AI choice, while affect and time discounting biases explain transitivity of AI choice over ballot formats.
 2.1 Fair collective choice is resilient to human-AI inconsistencies
-Figure 2 illustrates the human-AI consistency in individual and collective choices.
+Voting design and choice context have an impact on human-AI inconsistencies.
+Figure 2 illustrates the human-AI consistency in individual and collective choices for single-choice and multi-choice voting.
+For multi-choice voting, the individual and collective consistency of human and AI choices are measured as the average consistency across various sampled population sizes of 25%, 50%, and 75% (shown separately in Figure LABEL:fig:consistency-biases_actual and LABEL:fig:consistency-biases_survey).
 The consistency of individual choice remains poor in complex ballot formats with several alternatives.
-On average, it is 4.5% and 27.2% for the actual and mock voting in participatory budgeting, yet it is 82.3% for the binary majoritarian US elections.
-GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 5.2% and 5.31% higher than GPT 3 and Llama2.
-On the contrary, the consistency of the collective choice increases by 51.2% in total.
-Strikingly, the consistency of equal shares is on average 80.8%, which is 32.8% higher than utilitarian greedy.
-Even when reducing the number winners of equal shares to the 8 of utilitarian greedy, the consistency remains 29.4% higher.
-Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however consistency drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting.
-Strikingly, consistency of collective choice is significantly higher than individual choice, particularly for the fairer voting aggregation rule of equal shares.
-GPT 3.5 shows the highest choice consistency and Llama2 the lowest among the large language models, which though remain inferior to a predictive machine learning model.
-The consistency (y-axis) in individual and collective choice is shown for different AI models (x-axis) across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) mock, and (C) the US national elections of 2012, 2016 and 2020.
-The consistency in the participatory budgeting scenarios is shown for different voting input methods (left and right).
-The consistency of collective choice with the two voting aggregation methods of utilitarian greedy and equal shares is also shown.
-For the actual voting of City Idea, the consistency of equal shares is calculated for all winners and 8 winners (as many as greedy) for a fairer comparison.
-Compared to large language models, the machine learning model shows 3.6% higher consistency in individual choice and 6.6% higher in collective choice.
-All consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a.
-The removal of project preferences shows the highest consistency reduction of 19.7%, while political interest shows the lowest consistency reduction of 3.4%.
-Figure 3 illustrates the transitivity of preferences in different large language models by measuring the consistency of individual choice among different pairs of ballot formats.
-‘Approval - cumulative’ voting shows the highest consistency, which is 1.8% higher than ‘single choice - approval’ and 3.5% higher than ‘single choice - cumulative’.
-Figure 3: Voting personas based on GPT 3.5 show transitivity of the preferences over the different voting input methods in the actual complex voting scenario.
-However, GPT 3 and Llama2 remain with a low consistency <0.5.
-The simpler mock voting scenario shows higher transitivity among all language models.
+On average, it is 5.2% and 27.15% for the actual and survey voting scenario of City Idea, yet it is 82.3% for the binary majoritarian US elections.
+GPT 3.5 shows the highest consistency of individual choice among the large language models, which is 4.27% and 4.16% higher than GPT 3 and Llama3-8B.
+On the contrary, the consistency of collective choice increases by 45.03% in total.
+Strikingly, the consistency of equal shares and Phragmen’s is on average 68.8%, which is 30.7% higher than utilitarian greedy.
+Even when reducing the number winners of equal shares to that of utilitarian greedy, the consistency remains 21.4% higher.
+Compared to large language models, the machine learning model shows 3.3% higher consistency in individual choice and 4.2% higher in collective choice.
+The consistency values are the maximum ones derived with all personal human traits as shown in Figure 1a.
+Removing project preferences from the context of AI choice generation results in the highest consistency reduction of 18.1%, whereas political interest leads to the lowest reduction of 3.5%.
+Figure 2: Choice by large language models is consistent to humans for single-choice majoritarian elections, however accuracy drops for more complex ballots with larger number of alternatives as in the case of participatory budgeting.
+Strikingly, accuracy of collective choice is significantly higher than individual choice, particularly for the fairer ballot aggregation rules of equal shares and Phragmén’s.
+GPT 3.5 shows the highest consistency and Llama3-8B the lowest among the large language models, which though remain inferior to a predictive machine learning model.
+The mean consistency (y-axis) for different population of voters (10%, 25%, 75% and 100%) in individual and collective choice is shown for different AI models (x-axis), across three real-world voting scenarios: The participatory budgeting campaign of City Idea, (A) actual and (B) survey, as well as (C) the US national elections of 2012, 2016 and 2020.
+For participatory budgeting, the ballot formats of cumulative/score (left) and approval (right) are shown, including the ballot aggregation methods of equal shares, Phragmén’s and utilitarian greedy.
+For the actual voting of City Idea, the accuracy of equal shares is calculated for all winners and a controlled number of winners (as many as greedy) for a fairer comparison.
+Intransitivity: higher for AI with impact on collective choice among many alternatives.
+Figure 3 illustrates the transitivity of preferences in different large language models and humans by measuring the consistency of individual and collective choice among different pairs of ballot formats.
+While human transitivity approaches 97.1% on average, AI transitivity remains at 72.1% for GPT 3.5 and 78.3% for Llama3-8B.
+In terms of collective choice, equal shares has 12.2% higher consistency than utilitarian greedy.
+Equals shares has more than 80% consistency among winners from cumulative / score and approval ballots.
+For the actual voting of City Idea, ‘approval - cumulative’ voting shows the highest transitivity, which is 4.4% higher than ‘single choice - approval’ and 3.2% higher than ‘single choice - cumulative’.
+However, for the survey of City Idea, ‘single choice - cumulative’ shows the highest transitivity, which is 25.2% higher than ‘single choice - approval’ and 15.1% higher than ‘approval - cumulative’.
+Figure 3: Intransitivity of AI across different pairs of ballot formats is higher than the one of humans, which remains negligible.
+AI intransitivities have a higher influence on the consistency of voting outcomes over a large number of alternatives.
+Llama3-8B predicts ballots that are not very diverse and selects a limited set of projects, which results in higher transitivity compared to other language models.
+Equal shares and Phragmén’s also show here higher capacity to mitigate the ballot intransitivities.
+It achieves more than 80% consistency between cumulative and approval ballots.
 The consistency (y-axis) in individual choice among different pairs of ballot formats (x-axis) is shown for different large language models, humans and the two voting scenarios in the participatory budgeting campaign of City Idea: (A) actual vs.
-(B) mock.
-GPT 3.5 shows 21.1% higher consistency than GPT 3 and 20.2% higher than Llama2.
-While human transitivity approaches 95.2% on average, AI transitivity remains at 70.1% for GPT 3.5.
-The mock voting with only 5 alternatives achieves on average 65.5% of consistency, while the actual voting with 33 alternatives shows a consistency of 43.4%.
+(B) survey.
+Mean consistency values are calculated across randomly sampled population of 25%, 50%, and 75%.
 2.2 AI representatives to recover from low voters turnout
-Figure 4a illustrates the capability of AI representatives to recover the consistency of voting outcomes lost by low voters turnouts.
-Strikingly, up to 50% of AI representatives are sufficient for a maximum consistency recovery of 53.3% under voting turnouts lower than 80% when equal shares is used.
-This significant recovery is despite the very low consistency of human-AI individual choices in complex ballot formats with several options (Figure 2a).
-The recovery is also confirmed for equal shares with the top-8 winners, which is equal to the number of winners with utilitarian greedy (on average 18.1% lower consistency recovery).
-However, AI representatives fail to recover consistency with utilitarian greedy as the voting aggregation method.
-Figure 4: By representing up to half of the human voters who abstain from voting with AI, a significant lost consistency in the collective choice is recovered in participatory budgeting under fair voting aggregation methods such as equal shares and under low number of alternatives.
-However, further consistency loss is observed under a large number of alternatives with the standard utilitarian greedy method as the number of AI representatives increases.
-In contrast, single-choice majoritarian elections with highly consistent human-AI individual choice maximize consistency recovery for the whole spectrum of voters turnout by maximizing the AI representatives for the voters who abstain.
-Nevertheless, for more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI.
-The consistency loss in the voting outcome by low voters turnout (x-axis) is emulated by removing different ratios of human voters.
-A recovery of consistency (y-axis) is hypothesized by different ratios of abstained voters represented by AI: 25%, 50%, 75% and 100%.
-The consistency recovery values are the means of all three large language models (see Figure LABEL:fig:LLM_factorial for the results of each large language model).
-The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea are shown with two voting aggregation methods: utilitarian greedy and equal shares, with all winners and the top-8 (as many as greedy) for a fairer comparison.
-(C) US elections.
-(D) Chord visualizations summarizing how the interplay between voters turnout and AI representation yields voting outcomes with consistency gain or loss.
-(E) Visualizations summarizing the mechanism of consistency recovery through AI representatives for the voters who abstain.
-Figure 4b shows that the recovery of consistency is more robust for both voting aggregation methods under low number of alternatives.
-Consistency is recovered with up to 75% of abstained voters represented by AI under a maximum of 60% voters turnout, after which 50% of AI representatives are sufficient for consistency recovery.
-Under the voting scenario of the US elections that comes with highly consistent human-AI individual choices, consistency is fully recovered for $\geq$25of abstained voters represented by AI under a minimum of 50% voters turnout.
-However, below 50% of voters turnout, election results are very sensitive and consistency recovery is guaranteed only for 100% of abstained voters represented by AI.
-More detailed results on mixed populations of humans and AI representatives, including the consistency loss when human voters abstain are shown in Figure LABEL:fig:mixed_populations-abstaining.
-2.3 Biases explaining AI inconsistencies in choice and preference transitivity
-To explain the origin of (in)consistencies in human-AI choice and different voting input methods, a prediction model is designed for each voting scenario.
-The personal human traits, mapped to different cognitive biases, become the independent variables and the consistency the dependent variable.
-The prediction model is built using a neural network (see Section 4.4).
-The relative importance of the personal human traits that explain the AI consistency is calculated using shapley additive explanations (see Section 4.5).
-In the case of participatory budgeting in Figure 5a and 5b, the preference for environmental projects (potential conformity bias [40, 41, 42]) shows an average importance of 26.5% (p=0.0258) in explaining human-AI choice consistency (25% and 12% higher consistency among environmental supporters, with p=0.038 for the actual voting scenario).
-Political belief and ideology are strongly correlated with the actual two alternatives in the US elections (right/conservative for republicans and left/liberal for democrats), therefore, they strongly explain the consistency of human-AI choice (mean relative importance of 98.7%, p<0.02 and 4.8%, p<0.03 respectively).
-However, this is not the case in participatory budgeting as it is political belief that rather explains inconsistency (relative importance of -51.2% with p<0.02).
-This demonstrates that the underlying participatory budgeting choices remain politically neutral to a high extent, despite the fact that the City Idea campaign mainly attracted left-wing voters (42.5% versus 31.5% of right wing voters).
-In contrast, race and discussion of politics explain inconsistencies in the US elections with a mean relative importance of -48.1% (p<0.04) and -64.7% (p<0.04) respectively.
-Inconsistency is 33% higher than consistency among white people (p<0.019), while for people who discuss politics inconsistency is 63.2% higher that the ones who do not (p=0.046).
-Figure 5: Conformity bias, such as preference for environmental projects, positively contribute to the consistency of human-AI choice.
-On the other hand, unconscious biases such as race and discussion of politics cause inconsistencies.
-For US elections though, political belief and ideology strongly contribute to consistency.
-Time discounting factors such as preference for sport, and welfare projects as well as affect heuristics such as preference for projects that benefit families positively contribute to the consistency of large language models among the voting input methods.
-The relative importance (x-axis) of the personal human traits (y-axis) to predict the consistency of choice and consistency between voting input methods (single choice vs.
-cumulative/score) is shown for different large language models and the machine learning model.
-This is calculated using Shapley Additive Explanations.
-The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea is shown along with the (C) US elections.
-The ‘*’ denotes that the feature is statistically significant (p<0.05) for each of the AI models of Llama2, GPT 3, GPT 3.5 and ML respectively.
-The chord visualizations summarize the biases that explain the (in)consistencies of (D) human-AI choice and (E) voting input method.
-The consistency among the voting input methods is explained by preference towards time discounting factors [41] in participatory budgeting, such as sport and welfare projects.
-The average relative importance is 25.6% and 17.5% respectively (all with p<0.03), with 48.53%, and 37.5% of higher consistency among supporters in the actual voting (p<0.044).
-Affect heuristics such as preference for projects that benefit families with children also strongly contributes to the consistency with 58.9% higher consistency among supporters in the actual voting (p<0.032).
-Figures, LABEL:fig:merged3aLABEL:fig:explaination-top-choice, and Table LABEL:table:error illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.
+AI representation of abstaining voters is more effective than arbitrary voters under fair collective choice.
+Figure 4 illustrates the capability of AI representatives to recover the consistency of voting outcomes lost by low voters turnouts.
+The three abstaining models (low engagement, trust, digital literacy) are assessed along with the baseline that determines random abstaining voters across the whole population.
+We emulate all four participation modalities as shown in Figure 1a.
+Strikingly, up to 75% of AI representation of low-engaged abstaining voters (94 representatives out of 126 abstaining voters in a population of 252 voters, see Section LABEL:sec:abstainingStats) is sufficient to recover up to 50% higher lost consistency than the random control population using equal shares.
+This superior consistency recovery is also observed for abstaining voters with low digital literacy (39.96%) and trust (25.44%).
+The fair aggregation rules of equal shares and Phragmén’s achieve on average 7.53% higher recovery compared to greedy for all the abstaining models.
+Even when controlling for the same number of winners, fair ballot aggregation methods achieve higher recovery than greedy by 6.65%.
+On average, AI representation by GPT-3.5 achieves higher recovery across all abstaining models, outperforming Llama3-8B by 4.3% and GPT-3 by 6.1% (Figures LABEL:fig:consistency-recovery_remainLABEL:fig:consistency-recovery-survey-USLABEL:fig:consistency-recovery_llamaLABEL:fig:consistency-recovery_GPT3, and Table LABEL:tab:false_positives_negatives_all_models).
+Figure 4: Representing more than half of human abstaining voters with AI results in significant consistency recovery, in particular for fair ballot aggregation methods.
+Strikingly, AI representation of abstained voters is more effective than representing arbitrary voters (random control).
+Consistency recovery is at two levels: (i) False negative projects removed under abstaining but added back by AI representatives, which are are higher in ranking and number than (ii) false positive projects added under abstaining but removed by AI representatives.
+The consistency loss in voting outcomes by low voters turnout (x-axis) is emulated by removing different ratios of human voters (25%, 50%, 75% and 100%) among the whole population (baseline) and those who are likely to abstain: low engagement, trust and digital literacy profile (of the abstaining populations in the brackets on top).
+A consistency recovery (y-axis) is hypothesized by AI representation using GPT 3.5.
+(A) Actual participatory budgeting campaign of City Idea.
+(B) Studied participation modalities.
+(C)-(D) The origin of consistency recovery in participatory budgeting for utilitarian greedy and equal shares respectively.
+Abstaining voters result in falsely removing (left) and falsely adding (right) winning projects.
+AI representatives add back and remove these projects respectively to recover consistency.
+The projects and their probability to recover consistency under random control are shown for comparison.
+AI representation of arbitrary vs.
+abstaining voters: from removing noise to restoring representation deficit.
+Figure 4c and 4d show the origin of inconsistency under utilitarian greedy and equal shares when voters abstain and how AI representatives recover from this.
+The figures show which projects are involved in consistency recovery and their ranking: (i) erroneously removed projects (false negatives, left) that are correctly added back by AI representatives (true positive) and (ii) erroneously added projects (false positive, right) that are correctly removed by AI representatives (true negative).
+Compared to true negatives projects, true positive ones are higher in ranking by an average of 7.2 and 2.5 positions for equal shares and utilitarian greedy respectively.
+The higher consistency recovery by the abstaining models compared to the random control population originates from an average of 0.71 and 0.47 additional projects involved in consistency recovery for the two ballot aggregation methods respectively.
+Moreover, the origin of consistency recovery by abstaining models is more prominent to true positive projects (mean of 1.66 over 1.0 for true negatives), while it is more prominent to true negative projects in the random control populations (mean of 2.27 over 1.89 for true positives).
+See Table LABEL:tab:recovery for a complete outline based on all the AI models.
+This result demonstrates a distinguishing quality of targeting the AI representation to abstaining voters: representation deficit is restored by adding back winners who would not be there otherwise, while a non-targeted AI representation has a noise-removal effect by removing erroneous winners.
+The district wise consistency recovery for the Aarau has been enumerated in Table LABEL:table:region.
+2.3 Biases explaining AI (in)consistencies in choice and preference transitivity
+Unraveling biases that explain AI inconsistencies.
+Figure 5 illustrates the biases that explain the (in)consistency of human-AI choice and the AI transitivity among different ballot formats (single choice vs.
+cumulative).
+We mainly show the results of the actual participatory budgeting of City Idea, while the results of the other datasets are shown in Figure LABEL:fig:ML_supple.
+We distinguish between (i) the inconsistencies originated by the three identified subpopulation of abstained voters and (ii) the inconsistencies by the AI representation of the entire population.
+Prediction models are constructed using recurrent neural networks (see Sections LABEL:sec:bias and LABEL:sec:ml), demonstrating robust performance with F1 scores averaging over 80% for abstaining groups and 74% for the entire population across all large language models (Table LABEL:tab:result_all).
+The relative importance of the personal human traits (independent variables) that explain the AI consistency for individual voters (dependent variable) is calculated using model agnostic Shapley additive explanations and local interpretable model-agnostic explanations [43] (see Section 4.4, Figures LABEL:fig:Lime_approvalLABEL:fig:Lime_score).
+For the abstaining subpopulations, the dependent variable is the consistency difference (plotted in Figure LABEL:fig:consistency-gain) between the one of abstaining voters and the mean of 10 random control subpopulations as shown in Section 2.2 to better isolate the biases within these subpopulations.
+For the AI representation of the entire population, the dependent variable is the choice consistency.
+To provide more robust evidence, we distinguish in Figures 5c and 5d those personal human traits that explain AI consistency (i) in all datasets, (ii) AI models and (iii) those which are statistically significant (p<0.05).
+Figure 5: Compared to an arbitrary abstaining voter, those with low engagement and digital literacy exhibit characteristics that explain the consistency of human-AI representation and ballot formats, for instance, no interest in politics and support to education/health projects related to unconscious and surrogation biases.
+Time discounting, affect and conformity biases, such as preference for public space and environmental projects as well as support to families contribute to the consistency of human-AI choice.
+Time discounting factors such as preference for sport, and welfare projects as well as affect heuristics such as preference for projects that benefit families explain AI consistency among ballot formats.
+The relative importance of the personal human traits (y-axis) for the actual participatory budgeting campaign of City Idea over different large language models (x-axis) is depicted by the size of the bubbles and it is calculated using Shapley additive explanations.
+The consistency of human-AI representation (A) and ballot formats (B) (single choice vs.
+cumulative) is assessed.
+For each of these, the personal human traits explain the following: (i) The consistency difference between the three abstaining models and their random control.
+(ii) The (in)consistency of AI representation and transitivity for the whole population.
+The ‘-’ sign indicates non-significant values (p>0.05).
+(C)-(D) The statistically significant biases present in all AI models and datasets are summarized by chord diagrams.
+Voters likely to abstain explain the consistency of their AI representation and transitivity of ballot formats.
+No interest in politics, support to self-benefit and children explain the consistency in AI representation of abstaining voters with low engagement compared to a random abstaining voter (Figure 5c): mean importance of 16.13% (p=0.04), 16.81% (p=0.04) and 9.32% (p=0.03) respectively.
+These traits relate to potential unconscious, egotism and affect biases [44, 45].
+Potential surrogation and optimism biases stemming from support to education and public transit projects (mean importance of 17.23%, p=0.04 and 14.56%, p=0.04) explain the AI transitivity over different ballot formats for these abstaining voters (Figure 5d).
+The traits that explain the consistency in AI representation of abstaining voters with low digital literacy are support to city benefit (15.45%, p=0.02), children (8.11%, p=0.03), public transit (4.53%, p=0.04) and education (5.01%, p=0.02) that relate to altruism, conformity, optimism and surrogation biases (Figure 5c).
+Support to health (surrogation, 17.89%, p=0.03) and elderly (affect, 13.44%, p=0.03) projects explain the AI transitivity over different ballot formats for abstaining voters with low digital literacy (Figure 5d).
+These traits relate to older voters, who usually come with low digital literacy.
+Affect and unconscious biases explain the (in)consistency of human-AI choice, while affect and time discounting biases explain transitivity of AI choice over ballot formats.
+The consistency of human-AI choice is explained by support to families (affect, 7.43%, p=0.03), public space (time discounting, 8.91%, p=0.01) and environment (conformity, 8.46%, p=0.03), while inconsistency is explained by support to elderly (affect, 11.39%, p=0.04).
+For the US elections, a political profile of left explains consistency of human-AI choice, while white voters explain inconsistency (33% higher than consistency, p<0.019, see Figure LABEL:fig:ML_supple).
+Affect and time discounting biases also explain the transitivity of AI representation over different ballot formats, in particular the support to families (18.22%, p=0.01), welfare (16.78%, p=0.02) and sport projects (17.05%, p=0.02).
+Figures 5LABEL:fig:Lime_scoreLABEL:fig:approvalLABEL:fig:Lime_approval, Table LABEL:tab:approval_tab and Section LABEL:sec:ml illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.
 3 Discussion
-Generative AI voting: an inevitable race to safeguard democracy.
+The inevitable of generative AI voting and the race to safeguard democracy.
 Generative AI voting is likely to emerge as an inevitable technological convergence of AI and e-voting solutions currently adopted in real-world.
@@
 For instance, participatory budgeting elections are mainly conducted digitally, while Estonia has already institutionalized a digital identity for 99% of its citizens as well as electronic voting since 2005 [50].
-As Toomas H.
-Ilves, former president of Estonia, emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen".
-In the light of these converging advancements, the inter-operation of a generative personal voting assistant with a digital voting platform becomes technologically feasible along with the citizens’ need to have a more direct saying in several public matters and consultations.
-Therefore, the findings of this study become spot on to understand the implications of such a future, while they are significant to prepare safeguards for digital democracy.
-Fair voting design as a safeguard for generative AI voting.
+As the former president of Estonia emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen".
+In the light of these converging technological advancements, the inter-operation of a generative personal voting assistant with a digital voting platform becomes technologically feasible along with the citizens’ need to have a more direct saying in several public matters and consultations.
+Therefore, the findings of this study become spot on to understand the implications of such a future, while they are significant to prepare timely safeguards for digital democracy.
+Fair voting design as a democratic safeguard to generative AI voting.
 Large language models prove not ready yet to accurately represent humans in complex voting scenarios such as participatory budgeting.
+We disentangle a complex interplay of localized and diverse context of the project alternatives, the large number of those and the underlying multi-faceted biases to which large language models are susceptible.
 However, a striking finding of this study, built upon our demonstrated real-world evidence in Aarau, is that consistency of large language models becomes significant under fair voting aggregation methods that promote a proportional representation of voters’ preferences.
 This motivates a huge opportunity to get democracy "right" in the digital era of AI: move to alternative voting methods that yield fairer voting outcomes for all, while shielding democratic outcomes from AI biases and inconsistencies.
-The momentum for this win-win paradigm shift is tremendous, which would not be possible otherwise to demonstrate without the real-world empirical evidence based on which this study unfolds: fair voting rules such as the method of equal shares promotes under-represented projects and yield more winning projects with higher proportional representation of citizens’ preferences [39, 22].
-Nevertheless, how to scale up these democratic blueprints remains an open question, but AI can play a key role here.
+How to scale up these democratic blueprints remains though an open question.
+In particular, voting turnouts in participatory budgeting remain very low and far lower than other elections such as referenda or national elections.
+Despite eminent ethical and legal challenges of engaging AI representatives to democratic processes, our findings show that a more ethically-aligned AI representation of abstaining voters recovers consistency of voting outcomes, which would be lost in any case.
+This consistency loss can be to such large extent that is currently posing a long-standing barrier for participatory initiatives to take-off.
+Note that we focus on the AI representation of abstaining voters who intend to participate but their low engagement [41, 51], digital literacy [12, 52, 8, 53, 54] or trust [55, 56, 42] are a barrier for them.
+This is to distinguish voters whose abstaining is a conscious deliberate act and their AI representation would not be relevant or even desired in this context.
+Last but not least, our findings also demonstrate that AI representation alone does not suffice - a fair voting design is imperative to materialize significant recoveries from low voting turnouts.
 Why fair collective choice is resilient to AI biases and inconsistencies.
@@
 There is evidence that the method of equal shares has an inherent stability in the resulting voting outcomes [28].
-For instance, even with 25-50% of the voters, the elected projects do not significantly change [22].
-This is despite the fact that projects may end with very different number of total votes, as low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters.
+Low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters.
 Such projects are expected to be an origin of consistency.
-Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abstaining.
-As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study becomes the first of its kind and comes with compelling realism and merit for the validity of our significant findings.
+Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abs and Table LABEL:table:change_gr_ES that shows the origin of this stability in terms of new projects added and removed in the winning set.
+Nevertheless, equal shares is still affected by low voters turnout, especially, for participation rates $<50\%$ [28].
+As voters turnout in participatory budgeting is typically very low, these inconsistencies are both relevant and prevalent.
+Note that any comparison of stability between equal shares and utilitarian greedy should be made with caution as the number of winning projects under equal shares is much larger than utilitarian greedy.
+When we control for the number of winning projects between the two methods, equal shares remains more robust than utilitarian greedy, but to a much lower extent (see baseline [random] in Figure 4a and LABEL:fig:consistency-recovery_remain.)
+Real-world testing of equal shares: overcoming a validity barrier.
+As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study becomes the first of its kind with compelling realism and merit for the validity of our significant findings.
 This comes in stark contrast to other earlier studies [28, 27] that hypothesize the application of equal shares over proposed projects and ballots aggregated with the standard method of utilitarian greedy.
-Mitigating low voters turnout: AI representation is not enough, voting design matters.
-This perspective is assessed for the first time in this study by emulating voting scenarios with varying voters turnout, while letting voters who abstain to have a representation by generative AI to restore the lost consistency of abstaining.
-Nevertheless, the inherent inconsistencies of generative AI can hinder the effectiveness of AI representatives.
-Therefore understanding these opposing dynamics comes with significant merit to mitigate the impact of AI biases and inconsistencies in collective decision making.
-This is by showing striking evidence that voting design has a determining role on how much consistency can be recovered and what effect this can have for the legitimacy of the voting process and outcomes.
-Recovering the lost consistency under very low voting turnouts in single-choice majoritarian elections requires populations with very high majorities of AI representatives, which turns out to be illegitimate.
-In contrast, the lost consistency under higher voting turnouts can be restored even with low proportions of AI representatives.
-The fair method of equal shares is also here a prominent accelerator for the consistency recovery: fewer AI representatives are sufficient to restore consistency even under very low voter turnouts.
-It is also evident that such consistency recovery could not be achieved with the standard voting aggregation method.
-Trustworthy generative AI voting: a call for research and policy action.
+Towards a trustworthy generative AI voting: a call for research and policy action.
 What information large language models use to reason about voting decisions is influential for different types biases to manifest.
-Unconscious and conformity biases, related to political belief, race and environment reinforce human choices or yield inconsistencies.
-Time discounting biases, related to the support of projects on welfare and sport, explain preference transitivity among different ballot formats.
-These significant findings demonstrate that training data in generative AI voting are expected to be highly impactful for the representation of the voting outcomes in the population.
+This is particularly the case for affect, unconscious and time discounting biases involved in AI representation of human choices and the transitivity of AI choices over different ballot formats.
+Abstaining voters with a low engagement, digital literacy and trust also possess related personal human traits that explain the consistency of their AI representation.
+Voters who come with a more active participation profile, without typical features of abstaining voters, appear irreplaceable as the reasoning of large language models cannot accurately estimate their choices (Figure LABEL:fig:consistency-recovery_opp).
+This motivates a tailored, purposeful and finite use of AI representation with the aim to set itself obsolete by recovering participation of abstaining voters, while mitigating for the consistency loss as long as voters abstain.
+Training data in generative AI voting are expected to play a key role for representative voting outcomes of the voters population.
 Ethical and democratic guidelines are urgently needed, particularly for the use of (generative) AI in voting processes.
@@
 4.1 Emulating AI representatives
-We introduce a generalized design to create AI representatives using three large language models and a predictive AI method as a benchmark (referred to as ML) to generate choices (ballots) and emulate AI voting representation.
-Input to ballot generation.
+We introduce a generalized design to generate choices (ballots) and emulate AI voting representation using three large language models and a predictive AI method as a benchmark (referred to as ML).
+Input for ballot generation.
 This comprises of (i) defining the voting scenario, the voting alternatives and the personal human traits of the voter as the input context (LLMs) or features (ML) to inform the ballot choice, (ii) generating the ballots and (iii) aggregating the ballots into a collective decision (see Figure 1).
   • (i)
-    Personal human traits: Voter’s information is used to define an AI persona of a human voter, i.e.
+    Personal human traits: Voter’s information is used to define an AI persona to represent human voter, i.e.
 the AI representative (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2 in SI).
@@
 winning projects with proportional representation of citizens’ preferences).
-These traits are extracted from feedback surveys of voters, who also cast ballots in the voting scenarios and therefore the survey data are linked to the actual voting behavior (see field study design in Section 4.2).
+These traits are extracted from feedback surveys of voters, who also cast ballots in the voting scenarios and therefore the survey data are linked to the actual voting behavior (Tables LABEL:table:prologue1LABEL:table:prologue2LABEL:table:prologue3LABEL:table:prologue4LABEL:table:prologue5 in Section 4.2).
   • (ii)
@@
 They are illustrated in detail in Section 4.2.
-The actual City Idea scenario comes with 33 project alternatives and their cost, while the mock scenario comes with 5 costed project alternatives.
-The US elections comes with two alternatives.
+The actual City Idea scenario includes 33 project alternatives and their cost, while the survey scenario comes with 5 costed project alternatives.
+The US elections however is associated with two alternatives.
 These three scenarios provide a broad spectrum of decisions spaces to assess the AI representatives.
@@
 In the case of score ballots, the voters provide a score to a project from a specified range such as 1 to 5 with 1 being the lowest preference and 5 being the highest.
-In the case of cumulative ballots, the voters distribute a number of given points among the projects.
+In the case of cumulative ballots, the voters distribute a number of given points (i.e.
+10) among the projects.
 Single-choice ballots are derived from single approvals, top scores or top points distributed to a project.
 Generative AI.
-We use proprietary models such as GPT 3.5, GPT3 [50] and the open-source Llama2-13B [51].
-GPT 3.5 and GPT 4 have similar performance for these tasks [52], hence we resort to GPT 3.5 for the analysis.
-Furthermore, both GPT 3.5 and GPT 4 use the “Let’s think step by step” to evaluate a prompt and find a solution that mimics human reasoning [52].
-For each of the large language models, predictions are made based on the five combinations of personal human traits (see Figure 1), however the predictions with all models have the highest consistency to human choices, which we use for the analysis.
+We use proprietary models such as GPT 3.5, GPT 3 [59] and the open-source Llama3-8B [60].
+For each of the large language models, predictions are made based on the five combinations of personal human traits (see Figure 1b), however the predictions with all models have the highest consistency to human choices, which we use for the analysis.
+We explored advanced OpenAI models, including GPT-4 Turbo, for the survey scenario of the City Idea campaign across various input context combinations related to personal human traits.
+The responses obtained under identical hyperparameter settings closely resemble those of GPT-3.5, with mean score deviations of only $\pm$0.05 across all alternatives.
+Since both GPT-4 and GPT-3.5 employ the same prompt evaluation logic for reasoning tasks [61] for standard prompt engineering techniques, we opted for the more cost-effective GPT-3.5 model for all prompt context variations and the different datasets.
 Predictive AI benchmark.
 The personal human traits are set as features to predict the ballots using classification methods [62].
-We sample 40% of the data and use synthetic minority oversampling [54] to augment the data.
-This 40% is decided empirically based on the performance of the neural network models.
+Neural networks are used with 3 hidden layers and the relu [63] activation function for the sub-networks in the neural network.
+We employ the ensemble method [64] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [65].
+This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data (more details in Section LABEL:sec:ml).
+We sample 40% of the data and use synthetic minority oversampling [66] to create synthetic data for augmenting the dataset.
+We conclude that 40% is the optimal proportion based on a standard approach of conducting multiple experiments with varying ratios of synthetic and real data [67, 68] and analyzing their impact on model performance.
 We also divide the prediction into multiple sub-problems, which predict the approvals or assigned scores for every option and then compile the final ballot.
@@
 For the US elections, majority determines the winner.
-For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method or the method of equal shares.
+For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method [69] or the method of equal shares [39].
 Utilitarian greedy simply selects the next most popular project, the one with the most received votes, as long as the available budget is not exhausted.
@@
 It selects a project if it can be funded using the budget shares of those who voted for the project.
-A full explanation of the is out of the scope of this article and can be found in earlier work [38, 48, 49].
+A full explanation is out of the scope of this article and can be found in earlier work [39, 57, 58].
 In practice, the method of equal shares is likely to sacrifice an expensive popular project to select instead several low-cost projects, which, in combination, satisfy the preferences of more voters [27, 28].
 As a result, equal shares tends to elect more winning projects.
-Because of this effect, it likely that consistency measurements based on the Jaccard similarity yield higher values for equal shares.
+Because of this effect, it is likely that consistency measurements based on the pairwise similarity yield higher values for equal shares.
 This is the reason we control for the number of winning projects in equal shares by counting a subset of the most popular winning projects which is equal in number with the winners of the utilitarian greedy method.
-Emulated voting processes with human voters and AI representatives.
-We emulate the following voting processes: (i) Human voters exclusively with 100% turnout.
-(ii) Human voters exclusively with varying turnout in the range [10%,100%] with a step of 10%.
-(iii) Mixed populations of human voters and AI representatives assessed of whether they can recover the consistency of voting outcomes lost by different human voters turnouts in the range [10%,100%] with a step of 10%.
-The ratio of AI representatives for the abstained voters varies in the range [25%,100%] with a step of 25%.
-(iv) AI representatives exclusively.
-Each scenario of voters turnout is repeated 10 times by selecting a random subset of voters who abstain.
-For each turnout level, the AI representation of abstained voters is also repeated 10 times by selecting a random subset of the abstained voters to represent by AI.
+We also use the Phragmén’s proportional aggregation rule [40].
+Phragmén’s method balances fairness and representation between groups, whereas the Method of Equal Shares emphasizes fair representation within groups, ensuring that at least one voter from each group is represented [70].
+Modeling abstaining voters.
+We model three types of abstaining voters: (i) Those whose low digital skills may be a barrier to effectively use digital platforms to participate online and vote [12, 52], which often results in abstaining from participatory processes [8, 53, 54, 42].
+(ii) Those with lower political engagement [41, 51] and (iii) distrust to institutions [55, 56, 42] often choose to abstain from voting.
+We identify pre- and post-voting survey questions participatory budgeting campaign of City Idea (see Table LABEL:table:prologue3LABEL:table:prologue4 and LABEL:table:prologue5) that are proxies to determine potential voters with one of the three abstaining profiles [54, 71].
+We divide voters to two quartiles based on their responses to distinguish the voters that are likely to abstain and the ones that are likely to participate.
+This results in voters with low digital literacy, engagement and trust that characterize 48.25%, 36.12% and 27.36% of the population respectively.
+Emulating different participation modalities.
+Four participation modalities are studied: (i) Human voters exclusively with 100% of voters’ turnout.
+(ii) Human voters exclusively with varying turnout levels in the range [10%,100%] with a step of 25%.
+The maximum number of abstaining voters is either the total voters (baseline in Figure 4b) or the number of voters with low digital literacy, engagement and trust as determined by the abstaining models.
+(iii) Mixed populations of human voters and AI representatives of abstaining voters in the range [25%,100%] with a step of 25%.
+(iv) AI representations exclusively.
+A random control to rigorously assess AI representation of abstaining voters.
+The consistency loss by a number of abstaining voters belonging to one of the three groups of voters with low digital literacy, engagement and trust, as well as the recovery by their AI representation, is compared with the mean consistency loss and recovery of 30 groups of random voters, each of size equal to the abstaining voters (see Figure 4).
+For each turnout level of the abstaining voters, the process is also repeated 30 times by selecting a random subset of abstaining voters to represent by AI.
 Data collection infrastructure.
@@
 We prompted the large language models using the zero-shot learning feature [59], which does not require any specific fine-tuning.
-We use chain of thought prompting [55] to provide a comprehensive and systematic flow of information for better interpretability.
-The large language models run with temperature settings of 0.2 to 0 (20 runs for each temperature setting), for a focused decision with less randomness [56].
-As we are dealing with a significantly large decision space, particularly the 33 projects in the actual voting, running with a high-temperature setting can lead to random decision-making.
-To maintain consistency, we retain the projects that are approved at the deterministic temperature setting (the responses are easily reproducible at temperature setting = 0).
-Furthermore, if a project is selected both in 0.1 and 0.2, we include the approval for that project in the ballot to also incorporate diversity.
-In the case of score voting, we take the frequency of the distinct scores assigned to a project across different temperature setting and add the score with the highest frequency.
-However, for the actual City Idea scenario for which we emulate cumulative voting, the total score provided in a ballot should be lower than the total number of points.
-Therefore, it would not be be accurate to derive a score for a project from multiple outputs corresponding to different temperature settings, hence we conducted experiments only with the temperature setting of 0.2.
-The decoding algorithm for the large language model is set to the nucleus sampling approach [55].
-Compared to purely random sampling, it avoids generating outliers or improbable sequences of choices.
-For the predictive AI choices, we use neural networks with 3 hidden layers and the relu [57] activation function for the sub-networks in the neural network.
-We employ the ensemble method [58] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [59].
-This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data.
+We use chain of thought prompting [72] along with context-based prompting [73] to provide a comprehensive and systematic flow of information for better interpretability.
+A detailed explanation for the prompt designing is provided in Section LABEL:sec:prompt.
 4.2 Voting datasets
-Table 1 outlines the settings of the emulated voting scenarios.
+Table 1 outlines the characteristics of the emulated voting scenarios.
 US elections.
 The 2012, 2016, and 2020 survey waves of the American National Election Study (ANES) [74] are used.
-They encompass 4,500 votes together with the respective voter information for each of these three years.
-This mainly includes socio-demographic characteristics: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence.
-A total of 9 elections are emulated using Llama, GPT-3.5 and the predictive ML model based on single-choice ballots, with winners decided by majority for the three years.
+The dataset for three years contains 20,650 voters together with the respective voter socio-demographic characteristics for each of these three years: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence.
+A total of 18 elections are simulated using two combinations of human traits and three AI models, including Llama, GPT-3.5, and a predictive ML model based on single-choice ballots, with winners determined by majority vote over three years.
 We extracted information from 20,650 voters of the ANES dataset [74] and after pre-processing (removing voters with missing values in more than 5 of the personal traits), a set of 17,010 voters is used.
@@
 City Idea: Participatory budgeting campaign in Aarau, Switzerland.
-The data from a recent innovative participatory budgeting field study are used [32, 39], which was conducted with ethical approval from University of Fribourg (#2021-680).
+The data from a recent innovative participatory budgeting campaign are used [33, 27], which was conducted with ethical approval from University of Fribourg (#2021-680).
 It run in 2023 and is rigorously designed to assess the application of equal shares for the first time in real world, in combination with cumulative voting, using the open-source Stanford Participatory Budgeting platform [34].
@@
 We use the following personal human traits from the survey (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2): (i) 9 key socio-demographic characteristics (e.g., age, citizenship, education) and 2 political interests (political beliefs and trust in democracy), (ii) preferences for 9 different types of projects and 6 beneficiaries and (iii) 4 types of preferences for qualities of the voting outcome.
+Sc
 Table 1: The studied dimensions across three real-world voting scenarios.
-They provide the necessary diversity to generalize the findings of this study as they include different scales in number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases including both generative and predictive AI methods.
+They provide the necessary diversity to generalize the findings of this study as they include different number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases including both generative and predictive AI methods.
 Studied factors US Elections City Idea City Idea
-2012, 2016, 2020 [Mock] [Actual]
+2012, 2016, 2020 [Survey] [Actual]
 Ballot Input
 Personal human traits
-Socio-demographics
-Preference for voting outcome qualities
-Project preferences
-Political interests
+Socio-demographics, Outcome Expectations, Project Preferences, Political Interests
+Socio-demographics, Outcome Expectations, Project Preferences
+Socio-demographics, Project Preferences, Political Interests
+Socio-demographics, Outcome Expectations, Political Interests
+Socio-demographics, Political Interests
+Socio-demographics, Political Interests (only 1 feature)
 Ballot format
@@
 GPT 3.5
-Llama2-13B
+Llama3-8B
 Predictive AI
@@
 Equal shares
-Voters $\sim$17,000 (across 3 years) 3,134 505
-Emulated elections 9 91 27
+Voters $\sim$17,010 (across 3 years) 3,314 505
+Emulated elections 18 180 108
 Two participatory budgeting voting scenarios are studied in the context of the real-world campaign of City Idea:
   • (i)
-    Mock Voting: Five hypothetical costed projects belonging in different categories are put for mock voting as part of the initial survey.
+    Survey Voting: Five hypothetical costed projects belonging in different categories are put for choice as part of the initial survey.
 Table LABEL:table:prologueS illustrates the project alternatives and their cost.
-The choice of 3134 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting.
-This allows us to emulate a total of 45 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits + 1 for the predictive ML model.
+The choice of 3314 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting.
+This allows us to emulate 180 elections = 3 ballot formats x 4 AI models x 5 combinations of personal traits x 3 ballot aggregation methods.
+An additional 9 elections were emulated using all human traits, the three ballot format - (single choice, approval and score) and greedy and equal shares aggregation for the GPT 4, only to evaluate the differences with GPT 3.5.
+There were no difference in collective consistency and mean of the individual consistencies varies with $\pm$0.05 when compared to that of GPT 3.5.
   • (ii)
@@
 The ballot formats of single choice and approvals are derived from the cumulative ballots by taking the project with the most points and the projects that received any point respectively.
-This allows us to emulate a total of 91 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits.
-The examples of the prompts used to generate the AI choices in both mock and actual voting can be found in Table LABEL:table:prologuePersona.
+This allows us to emulate 108 elections = 3 ballot formats x 4 AI models x 3 combinations of personal traits x 3 ballot aggregation methods.
+The examples of the prompts used to generate the AI choices in both survey and actual voting can be found in Table LABEL:table:prologuePersona.
 4.3 Evaluation of choices by AI representatives
@@
 the resulting voting outcomes.
-Consistency is the main measure of assessment.
-Consistency of individual choice. Consistency is assessed at two levels:
+Consistency is the key assessment measure, derived from the accuracy of individual and collective AI choices compared to human decisions and the transitivity across different ballot formats (see Figure 1a).
+Consistency of individual choice.
+Single-choice ballots for both AI and human voters are represented as binary sequences, where a value of 1 indicates approval of a specific project, and 0 denotes disapproval of all remaining alternatives.
+In approval voting, each alternative is assigned either 1 (approved) or 0 (not approved).
+In contrast, in score voting and cumulative voting each alternative receives a score or an number of distributed points (integer numbers) reflecting voter preference.
+To compare AI-generated and human choices, we employ a single method, the Condorcet pairwise comparison method [22, 13], which is a generic approach to characterize the overall similarity of two ballots (or voting outcomes).
+A preference matrix is constructed, where rows and columns correspond to alternatives, and each matrix element records the outcome of a pairwise comparison.
+If project $P_{i}$ is ranked higher than project $P_{j}$, or if $P_{i}$ is approved while $P_{j}$ is not, the corresponding matrix cell $P_{i}>P_{j}$ is assigned a value of 1; otherwise, it is set to 0.
+Ties are excluded from the analysis.
   • (i)
-    Human-AI consistency of individual choice: This is the Jaccard similarity [61] of individual choices made by AI representatives with the original choices of the humans they represent.
-For single-choice ballots, this is either 1 (consistent) or 0 (inconsistent) for each voter and the respective AI representative.
-For approval voting, the AI and human choices are binary sequences representing the ballot, containing 1 (approved) or 0 (not approved) for each alternative.
-The Jaccard similarity between these sequences is the ratio between the number of alternatives approved by both the human and the AI representative over the number of alternatives approved by at least one of them.
-For score and cumulative voting, projects are ranked based on the scores and points received respectively.
-Each top-k project of a human voter is compared to the respective top-k project by its AI representative.
-Matching is indicated by 1 (consistent), otherwise 0 (inconsistent) indicates no matching.
-The mean number of matchings among all ranked projects measures the consistency for these ballot formats.
+    Human-AI consistency (accuracy) of individual choices: The human ballots serve as the reference point for evaluating the ones generated by the AI representatives, see Figure 1a.
+The elements of ‘1’ in the matrix of AI representatives that match the elements of ‘1’ in the matrix of human choices determine the consistency [13].
   • (ii)
-[…]
+    Consistency (transitivity) of AI and human individual choice across ballot formats: Ballot formats are standardized as follows: For cumulative/score vs.
+single-choice ballots, the highest-scoring projects are set to ‘1’ and the others to ‘0’.
+For cumulative/score vs.
+approval ballots, scored projects are set to ‘1’, while projects without score are set to ‘0’.
+The elements of ‘1’ and ‘0’ in the two matrices of the ballot formats that match determine the consistency.
+We also compare the choices based on preference reordering using the Kemeny distance [75] as illustrated in Section LABEL:sec:supllemethod.
 Consistency of collective choice.
-It is defined for two voting processes that elect winners from the same alternatives, but have different voters (human, AI representatives, mixed), and, as a result, can yield different voting outcomes.
-The voting outcomes with the winners are first calculated using one of the ballot aggregation methods, i.e.
-majority, utilitarian greedy or equal shares.
-These outcomes are represented as binary sequences containing either 0 (not winner) or 1 (winner) for each alternative.
-Consistency is calculated using the Jaccard similarity [61] between these binary sequences.
+This follows the same approach of Condorcet pairwise comparisons for individual choices.
+However, before calculations of consistency are made, voting outcomes are turned into binary sequences to distinguish winners (‘1’) from losers (‘0’) as determined by a ballot aggregation method.
 Consistency recovery in collective choice with AI representatives.
-It is determined here for voting processes with varying voters turnout, in which abstained voters result in consistency loss, which can be recovered if a portion of these abstained voters are represented by AI.
-It is measured as follows:
-$\frac{consistency\ [\text{human\ voters - abstained\ voters + AI\ representatives}]-consistency\ [\text{human\ voters - abstained\ voters}]}{1-consistency\ [\text{human\ voters - abstained\ voters}]},$
-where the voters turnout human voters - abstained votershuman voters + abstained $\frac{\text{human\ voters - abstained\ voters}}{\text{human\ voters + abstained\ voters}}$ varies in the range [10%,90%] with a step of 10%, and AI representation AI representativesabstained $\frac{\text{AI\ representatives}}{\text{abstained\ voters}}$ varies in the range [25%,100%] with a step of 25%.
-4.4 Human cognitive biases in AI collective decision making
-Human choices are significantly influenced by potential cognitive biases that are often a manifestation of socio-economic characteristics, conditions of life quality, satisfaction with the available public amenities and the overall life experiences of an individual [41].
-Here we map self-reported personal human traits with which large language models are prompt to potential underlying human cognitive biases, as a way to explore whether they are reinforced by the language models and become more likely to manifest in the AI choices.
-Figure 1c outlines the mapping we study based on a systematic and comprehensive review of relevant literature.
-The following types of biases are determined:
-Time-discounting biases.
-These are characterized by the tendency to receive immediate gratification over a larger but future reward.
-Projects related to the public space, culture, or welfare are often one-off projects, such as an annual festival, a cinema night, or educating asylum children (alternatives proposed for the participatory budgeting campaign in Aarau [32]), where expected benefits are experienced in the short term.
-Therefore, such project are subject of time-discounting biases [41].
-In contrast, the construction of a road, for example, incurs investment costs several years before the actual project implementation.
-Such investments create a sustainable reform that benefits all in the long run.
-Even though 72% of the public transportation infrastructure in European cities ends up in cost overruns, people still vote for these projects due to an underlying necessary optimism to improve transport [41, 39].
-We refer to this economic discounting over a long-term sustainability impact as ‘optimism’ [62, 63].
-Surrogation biases.
-This reflects how humans favor simpler measures to assess the impact over ones that are more precise and harder to evaluate.
-Korteling et al.
-[41] argue that these biases manifest when deciding projects with large societal impact such as health or education, while their outcome is subject of different satisfaction levels among citizens.
-The project outcomes may be perceived as successful by part of society using easy-to-evaluate metrics instead of looking at long-term effects on the community.
-For instance, a timely vaccination drive may be preferred over significant changes to vaccination protocols or health insurance policies covering vaccination.
-Hence these projects are likely to be more preferred as they come with more intuitive ways to assess for the broader population.
-This is reflected by the average winning rate of the health and education-related projects of 38.1% and 36.2% respectively in Poland [39], where participatory campaigns have been actively hosted in the last decade.
-Conformity biases.
-These biases arise out of group pressure under which people make decisions to be socially desirable [40, 42].
-It is argued that a conformity bias may induce voting for green alternatives [41].
-Green-themed participatory budgeting campaigns have been adopted in European cities such as Lisbon [64], to promote green initiatives, aligning to a culture for more sustainable behavior.
-Poland runs participatory budgeting campaigns at large scale, which include environment and urban greenery projects.
-These are within the top-5 most popular projects with an average of 22.5% and 26.5% respectively [39].
-Even in Aarau we observe the same trend wherein, environmental friendly projects account for the top-10 most popular projects [32].
-Affect heuristic biases.
-This defines the tendency to make decisions based on what intuitively or emotionally feels right.
-Affect biases have been studied to analyze the inclusive attitudes most people show towards elderly people [65].
-Similar biases also manifest in welfare of children and in inter-generational communication [41, 66].
-In Aarau, we observe that 71.3% of the voters prefer projects for younger and elderly people.
-Biases for altruism and egotism.
-Individual interest is often in conflict with the community interest in participatory and collective decision-making processes.
-While voting, the intrinsic altruism of citizens strongly influences choices.
-As a result, altruism and egotism are influential for the fairness of voting outcomes and how these outcomes benefit the city in overall [67, 41, 66].
-In Aarau, we observe that 67.1% of the voters, who prefer better representation in the outcome are prosocial and prioritize the city-wide benefit (altruism bias) over individual benefit (egotism bias).
-Unconscious biases.
-Human choices are influenced by socio-economic and demographic traits such as race, ethnicity, citizenship, household size and income [68].
-Specifically, political ideology and belief shape to a high degree decisions for candidates in elections [69].
+It is determined here for voting scenarios with varying voters turnout, in which abstained voters result in collective consistency loss, which can be recovered if a portion of these abstained voters are represented by AI.
+This recovery takes place at two levels: (i) False negative projects that are erroneously removed under abstaining but added back by AI representatives.
+(ii) False positive projects that are erroneously added under abstaining but correctly removed by AI representatives.
+Consistency recovery is measured as follows:
+$\frac{consistency\ [\text{all\ human\ voters - abstained\ voters + AI\ representatives}]-consistency\ [\text{all\ human\ voters - abstained\ voters}]}{1-consistency\ [\text{all\ human\ voters - abstained\ voters}]},$
+where the voters turnout human voters - abstained votershuman voters + abstained $\frac{\text{human\ voters - abstained\ voters}}{\text{human\ voters + abstained\ voters}}$ varies in the range [20%,75%] with a step of 25%, and AI representation AI representativesabstained $\frac{\text{AI\ representatives}}{\text{abstained\ voters}}$ varies in the range [25%,100%] with a step of 25%.
 4.4 Explainability of generative AI voting
-The consistency of the individual AI choices (see Section 4.3) is modeled as the dependent variable in a predictive machine learning framework.
-Causal relationships that explain how personal human traits (independent variables) contribute to consistency are studied.
-We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$.
-Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set.
-We also assess prediction models with different dependent variables, see Figure LABEL:fig:explaination-top-choice for more details.
+The accuracy of the individual AI choices (see Section LABEL:sec:explain) with human choices as well as the transitivity of AI choices over different ballot formats are modeled as the dependent variable in a predictive machine learning framework.
+We study causal relationships explaining how personal human traits (independent variables) influence consistency (both accuracy and transitivity).
+We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as the ranges $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$.
+Further details about how we account for imbalances of features, their colinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set.
 Explainability of choices.
-We introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on the consistency of individual choice.
-We use model agnostic Shapley Additive Explanations [70] to extract individual contributions of each trait.
-Results are shown in Figure 5.
-Feature ablation study [71] is used to calculate the error (loss) in the overall prediction accuracy of the model when a feature is removed (results in Table LABEL:table:error).
+We introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on the consistency of individual choices.
+For a given performance of the prediction model, we employ explainable AI methods to analyze the contribution of each individual human trait (feature) to the outcome.
+The approach to enhance the performance (accuracy, precision, recall) of the prediction model is illustrated in Section LABEL:sec:ml.
+We then use the model agnostic Shapley Additive Explanations (SHAP) and Local Interpretable Model Agnostic Explanations (LIME) [43] to extract the individual contributions of each trait.
+Results are shown in Figure LABEL:fig:ML_suppleLABEL:fig:Lime_scoreLABEL:fig:Lime_approvalLABEL:fig:approval and Table LABEL:table:error.
+A feature ablation study [76] is used to calculate the error (loss) in the overall prediction accuracy of the model when a feature is removed (results in Table LABEL:table:error).
 Acknowledgements